### v1

Abstract
The neuronal electric operation can entirely be described by physics provided that we take into account that slow positive ions instead of fast light electrons are the charge carriers that biology uses. Based on the slow ions, we derive the correct physical model that Hodgkin and Huxley wanted. By using laws of electricity and thermodynamics in a non-disciplinary way, we derive the resting potential from the first principles of science, from the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions). We show that a simple control circuit regulates neuronal operation, between a resting and a transient state and that issuing an Action Potential is part of the control process. We explain that the resting potential is the consequence of charge separation instead of an ad-hoc linear combination of mobilities or reversal potentials as the Goldman-Hodgkin-Katz equation claimed. We defy the hypothesis suggested by Hodgkin and Huxley (HH) that some "leakage current" produces the resting potential; this way solve the decades-old mystery of "heat absorption" by neurons, the controversy of the experienced and theoretically expected energy consumption of neural computing. We fix the mistakes HH guessed in the physical picture behind their empirical model and mathematical equations, and introduce the correct physical model by providing also its detailed mathematical description.
keywords:
slow ions, neuron’s physical model, resting potential, leakage current, negative feedback, thermodynamic driving force, Goldman-Hodgkin-Katz equation, irreversible heat production
1 Introduction
The excellent textbook [1] separates the neuron membrane’s states into resting and transient states. Concerning the resting state, the statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; here we tell the second half. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential"  [1], page 126. In the resting state, two effects are considered. On the one side, the potential is created by a static charge-up: "the cell membrane potential results from the separation of net positive and net negative charges on either side of the membrane". Furthermore, as we discussed, it is only a small part of the potential: the polarization of the electrolyte’s dipoles contributes to the rest. (For the sake of clarity: both mechanisms are known in physics. The word ’polarization" is reserved for rearranging charges in multiply-charged systems without external current. The changes in potential difference is the consequence of complete charge separation, even if physiology calls it, for example, hyper-polarization and repolarization. External currents are associated to the changes of potential.) On the other hand, it was claimed that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [1], although it was question marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [2]. The claim is rooted in the hypothesis and the measurements by Hodgkin and Huxley [3], as we discuss in section 3.5. Its numeric figure was correct, but it was mistakenly interpreted as being produced by a "leakage current" on the membrane’s parallel resistance. Paper [4] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways but did not account for the electric potential. We derive how that electric potential is produced by separating charges and how its value can be calculated from the first principles.
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water." [5]. More precisely, inseparable thermodynamic and electrical processes keep the potential set by the inseparable electric and thermodynamic backbone defined by the lipid structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [6].
The book asks the central questions, "How do ionic gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, as it discusses only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization.
We demonstrate that classical electric methods enable us to calculate the membrane’s potential resulting from charge separation and polarization. However, some physically plausible assumptions must be made to provide numerical figures. The electric processes are inseparable from the corresponding thermodynamical processes [7], which we discuss by deriving a "thermodynamical electrical field". The other minor effect is a low-intensity "passive flux" (material transport) produced by the "resting channels". They are ion channels not necessarily only across the membrane; they also include the ion channels in the Axon Initial Segment (AIS) [8, 9, 10, 11, 12]. The "resting current" is about two orders of magnitude lower [13] than it was assumed at the time of writing the book; correspondingly, its role is much smaller in the transient state, producing an Action Potential (AP) which we discuss elsewhere [14].
The importance of the subject was correctly estimated decades ago [15]: "For the past forty years our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of Hodgkin and Huxley. Their work has had far reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. These include not only neurophysiology but also endocrinology, muscle and cardiac physiology, and developmental biology. The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system. But ionic fluxes also play important roles in affecting cellular functions such as secretion, contraction, migration, etc." To fix the flawed hypothesis has a similar importance.
2 Membrane’s electricity
It was already guessed that the membrane’s resting potential is of electrostatic origin [16, 17, 18, 2, 19]. When ions are confined in a closed volume, they exist in a state of thermal and electrical equilibrium. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [7]. Furthermore, the features of living matter may change when measured in physics laboratories as usual. When measuring them, we must consider that "the construction is different from anything we have yet tested in the physical laboratory" [20].
In the absence of external influences (that includes the lack of a separating membrane), both gradients are balanced and are at zero. When this state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [7], and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions leads to incorrect conclusions.
2.1 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and thermodynamic environments. A molecule has internal electric forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electric field (which can be of electrical or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electric field fluctuates, so their state is dynamic: the molecules dissociate, and the free ions recombine. Depending on their environment, the substances can exist in base, ionized, and polarized states (and behaviors).
In biological cells, a small (approximately $10^{-3}$) portion of the molecules dissociates (i.e., the ions separate from their counterparts with opposite charges), and they move freely within the volume. In other words, the electrolyte liquid can then conduct electricity due to the mobility of the positive and negative ions. The rest of the molecules can be more or less polarized, allowing for the possibility of producing an internal electric field in the solution.
2.2 Deriving "thermal electric field"
In a segmented electrolyte, the electric charges are typically globally balanced; however, they may become unbalanced locally due to physical reasons. The Nernst-Planck electrodiffussion equation in a balanced state
$\frac{d}{dz}V_{m}(z)=-\frac{RT}{q*F}\frac{1}{C_{k}(z)}\frac{d}{dz}C_{k}(z)$ (1)
describes the mutual dependence of the spatial gradients of the electric and thermodynamic fields on each other. In good textbooks (see, for example, [21], Eq (11.28)), its derivation is exhaustively detailed. In the equation, $z$ is the spatial variable across the direction of the changed invasion parameter, $R$ is the gas constant, $F$ is the Faraday’s constant, $T$ is the temperature, $q$ the valence of the ion, ${V_{m}(z)}$ the potential, and ${C_{k}(z)}$ the concentration of the chemical ion. The two sides are the derivatives of the equation
$V_{m}=\frac{RT}{q*F}\ln{\biggl{(}\frac{C_{k}^{ext}}{C_{k}^{int}}\biggr{)}}$ (2)
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration). Hence, the derived potential inherently comprises an additive term (a potential difference), so they are not intercomparable directly if they use a different reference $C_{k}^{int}$. The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [7].
(It is a common fallacy in physiology that the applied potential $U$ generates a current. Instead, its gradient ($E=U/d$) has the effect, where $d$ is the distance, for example, between the clamping point and the membrane or the AIS.)
We assume that, in a balanced state, the concentrations on the two sides of the membrane are $C_{k}^{ext}$ and $C_{k}^{in}$, respectively; furthermore we assume that
$\frac{dC_{k}}{dz}\approx\frac{C_{k}^{ext}-C_{k}^{in}}{d}$
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, and that $U=E*d$, we can derive the equivalent "thermodynamic electric field" for the case of a permeable ion channel in the membrane
$E_{thermal}^{C_{k}}=-\frac{RT}{q*F}\frac{1}{C_{k}^{ext}}\frac{C_{k}^{ext}-C_{k}^{in}}{d}\approx-\frac{RT}{q*F}\frac{1}{d}$ (3)
It does not depend on the concentrations and depends reciprocally on the membrane thickness (given in $nm$); furthermore, it is valid separately for all $C_{k}$
$E_{thermal}^{C_{k}}(d)=-2.585*10^{-2}\frac{1}{d}\quad\biggl{[}\frac{V}{m}\biggr{]}$ (4)
For a $5\ nm$ membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Equ.(18). Figure 1 depicts its dependence on the membrane’s width and how it compares to the charge-generated electric field. Correspondingly,
$U_{thermal}^{C_{k}}(5)=-25.85\quad\biggl{[}{mV}\biggr{]}$ (5)
Figure 1: The thermodynamic "electric field" depends on the membrane’s thickness, see Eq.(4). The electric gradient depends on the charged layer’s concentration and thickness; see Eq.(18). The dots mark measured data shown in Table 1.
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electric field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, a joint charge-generated electric field exists, and they equal, per ion, the concentration-generated "electric fields", see section 3.2.1.
In the steady state, some other forces also contribute to the mentioned forces. To discuss how nature restores the steady state when a microscopic change happens in a balanced state of a biological solution, we write the well-known Nernst-Planck equation (see Eq.(1)) in a slightly extended form:
$0=\underbrace{e_{el}*E_{Gap}^{Total}(c,\Delta z)}_{Electrical~{}force}+\underbrace{e_{el}*E_{thermal}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{k}}}(d)}_{Thermodynamical~{}force}\quad\underbrace{\bigl{(}+F_{ext}(z)\bigr{)}}_{Constraint}\quad\underbrace{\bigl{(}+F_{Transp}(z)\bigr{)}}_{Transport~{}force}$ (6)
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added an external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid.
Expressed differently
[…]
That means when describing an ionic transfer process, we must not separate the electric current from the mass transfer: they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electric term is not. To be entirely balanced, the system must be balanced to all elements. In this way, changing one concentration implicitly changes all other concentrations and the electric field.
It is a colossal mistake to introduce equivalent circuits with their fixed-value voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason and prevents understanding how the competition between thermodynamic and electric processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 2.6; in this way attributing the change of an electric entity to the biological material. This assumption neglects the driving force the mentioned forces provide. Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
When discussing balanced states, no transport occurs, so the transport force cancels, and the mobility has no role. The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions, and the membrane mechanically provides the needed counterforce. If the boundary of the segments is not freely penetrable, the counterforce equals the difference between those two forces. In this way, no force acts on the ions; the two gradients persist. When the ions can move freely between the segments, they will move until they produce a concentration gradient (a thermodynamical force) for the given ion that counterbalances the electrical gradient (the electrical force) and then the transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq.(27)) in a viscous fluid.
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes the ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electric current in solids.
2.3 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line, perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electric effects connected to membranes and nanopores [17] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
In the calculations below, we require the ’surface charge density’ (interpreted as an ’infinitely thin layer’ in physics, expressed in $C*m^{-2}$). We can derive (assuming singly-charged ions) the volume charge density
$\sigma_{V}(c)=c*N_{A}*e_{el}=c*6.023*10^{20}*1.602*10^{-19}=96.4*c\quad\biggl{[}\frac{C}{mM*m^{3}}\biggr{]}$ (8)
where $N_{A}$ is the number of ions in a $millimol$ and $e_{el}$ is the elementary charge. Notice that all ions contribute to the charge density, so simply $c=\sum_{k}C_{k}$.
Here, we run into a conflict between the ’infinitely thin’ layer of physics and the biologically implemented (finitely) ’thin physical layer’, so we use a thickness parameter $\Delta z$. By assuming an arbitrary ’physical layer thickness’ $\Delta z$, we can calculate the surface charge density
$\sigma_{A}(c,\Delta z)=96.5*c*\Delta z\quad\biggl{[}\frac{C}{m^{2}}\frac{1}{mM*nm}\biggr{]}$ (9)
We interpret classical physics’ notions from the theory of ’continuous’ electricity, interpreted for ’infinite’ cases, for the finite world of biological objects and the quantized atomic world. Our idea is similar to the one used by Boltzmann for connecting the continuous medium to quantized particles. When ions are present in an electrolyte having concentration $c$ in a ’conducting layer’ of thickness $\Delta z$ on an isolating surface, they produce a field
$E_{z}(c,\Delta z)=\frac{\sigma_{A}(c,\Delta z)}{\epsilon_{o}}=10.89*10^{3}*c*\Delta z\quad\biggl{[}\frac{V}{m}\frac{1}{nm*mM}\biggr{]}$ (10)
Figure 2: The electric field of a disc of charge can be found by superposing the point charge fields of infinitesimal charge elements. This field can be facilitated by summing the fields of charged rings. “From Hyperphysics by Rod Nave, Georgia State University". ©hyperphysics.phy-astr.gsu.edu/hbase/
2.4 Two segments
Now, we separate the volume into two segments and compose the segments from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electric field’s $z$ component of those plates in the points as shown in Fig. 2. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electric field is zero inside the conducting layer), so the electric field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into the ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [22]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electric fields. Although not explicitly stated, we consider electric double layers to be present near the membrane, and ions from the dispersion medium are adsorbed on the particle’s surface, depending on the chemical properties of the medium.
Correspondingly, the theory non-equivocally explains the potential near the membrane surface; see the discussion and review in [23], mainly their Figure 1. That model considers a "symmetric" condenser, that is, both plates contain charges of the same sign. The Donnan function is a (discontinuous) step-like potential. Its continuous version introduces a mathematically smooth, non-physical function at the boundary of the charged layer. One must describe a physically discontinuous layer (electrolyte, ion layer on the membrane’s surface, and isolating membrane) with a single continuous function and gradients (potential and concentration) with physical meaning.
The behavior near the boundary layers is problematic, mainly because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electric field; the charged layer and the electrolyte jointly produce the electric field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electric field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their electric force is a fragment of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it creates an electric field similar to a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electric field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separates the segments). The membrane does not need to generate any counterforce (except for collisions).
2.4.1 Charge layers in segments
As we show, the finite thickness will result in a lack of balance (introduces inhomogeneity and creates a voltage and concentration gradient) proximal to the surfaces of the membrane, even if the concentrations on the two sides are the same. Changing the bulk concentration or potential in one of the segments creates a corresponding gradient across the separating membrane, increasing the inhomogeneity proximal to the membrane. The ions will experience an extra force due to the gradient; however, the mechanical counterforce of the membrane will keep them back in the segment. The concentration and potential, inseparably and having the same time course, will change across the two sides of the membrane, just because of the gap’s physical features the membrane represents.
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface) and the charged discs’ contribution to the electric field at point $P$ (see Fig.  2) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electric field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
Notice that here we used the abstraction of an infinitely thin "charged sheet", and there is a step-like gradient in the electric field in the gap. The physical reality is that the charge is represented by ions, which simultaneously represent a mass. In equilibrium, the forces due to the voltage and concentration gradients must be counterbalanced by an external force. An interference between science disciplines can also manifest here. We know simultaneously that at the boundaries of electrolytes, different interfaces, including electrically neutral electric double layers, can be formed by only partly known processes [24]. The presence of those structures makes drawing quantitative conclusions hard.
2.4.2 Simple conducting sheet (a physical condenser)
First, we consider a condenser that obeys the laws of classical electricity and has the geometrical size of a biological neuron. We apply the physics terms to that pseudo-biological object and test whether the values we derive are consistent with physiological phenomena. A tiny (according to [25], page 12, about $2*10^{-3}$) portion of the positive ions leaves their negative counterions behind and forms a thin positively charged ion layer on the surface of the membrane (of course, the same happens on the opposite side with negative ions). Experience shows that – also in biology – two very thin layers of ions are formed on the two sides of the membrane of finite width, see the caption of figure (Fig. 11.22 in [26]): "The ions that give rise to the membrane potential lie in a thin ($<1\ nm$) surface layer close to the membrane" (As shown below, these ions give rise only to part of the potential, and the dipoles in the electrolyte contribute the rest). The two surfaces are covered by a sub-nanometer-thick conducting layer on the top of a $5~{}nm$ thick insulator. So, we model the cell with two finite-width "conductive sheet" layers. These layers represent an electric condenser (so we can calculate the internal electric field between the plates) and counterbalance each other’s electric field outside the condenser. We may assume that the solid surface represents the needed counterforce to keep the charges at rest.
We cannot expect geometrically flat surfaces, as some biological objects of up to $1~{}nm$ in size are present on the surface. So, we must assume a physical ’uniformly charged sheet’ with an average thickness $\Delta z=0.5~{}nm$ on the surface. For the sake of simplicity, we assume a step-like function for the electric field. It is zero in the segment outside the "conducting sheets" (the ’bulk’ portions) and inside the sheets. Between the sheets, it jumps to the value of the electric field of a charged sheet with surface density $\sigma_{A}$ (see Eq.(9)). (In this picture, we see a jump in the value of the electric field on the two sides of the physical layer. We assume that the physical layer somehow emulates the well-defined boundary on the side opposite to that proximal to the membrane. Here, we still assume that a mechanical counterforce due to the surface’s roughness keeps the charges inside the charged layer.)
We assume that all unbalanced (dissociated) ions are in the layer, and the rest contains no dissociated ions; furthermore, the concentration of the ions is unchanged in the segment. In the case of a $400~{}mM$ solution and a $0.5~{}nm$ ’thin physical layer’ Eq.(10) evaluates to an electric field
$E^{Gap}_{Classic}(400,0.5)=0.22*10^{7}\biggl{[}\frac{V}{m}\biggr{]}$
and at a $5\ nm$ membrane thickness, the voltage across the plates becomes
$U^{Gap}_{Classic}(400,0.5,5)=E^{Gap}_{Classic}(400,0.5)*10^{-3}*d=10.84\ \bigl{[}mV\bigr{]}$ (11)
Our simple model seems to be a strong oversimplification. The number of uncompensated ions needed for the cell, using the method in [25], is $0.88*10^{7}$. The calculated values are about a factor of $5$ lower than the experimentally derived values. "An electrical potential difference about $50-100\ mV$ exists across a plasma membrane only about $5\ nm$ thick, so that the resulting voltage gradient is about $100,000\ V/cm$[26]. "The number of uncompensated ions needed for the cell is $4.7*10^{7}$[25]. We are in the correct order of magnitude, but we arbitrarily assumed a layer thickness, a non-dielectric medium, and a physically unreasonable step-like electric field. The deviation from the experienced values suggests that dipoles in the electrolyte bulk significantly change the achievable electric field and potential across the plates.
Figure 3: The oppositely charged parallel surfaces of the membrane are treated like conducting plates (infinite planes, neglecting fringing), then Gauss’ law can be used to calculate the electric field between the plates. Presuming the plates are at equilibrium with zero electric fields inside the conductors, the result from a charged conducting surface can be used. “from Hyperphysics by Rod Nave, Georgia State University". ©http://hyperphysics.phy-astr.gsu.edu/hbase/
A usual parallel plate condenser (shown in Fig. 3) can be derived with such a model. We have two infinite conducting sheets and an insulator layer between them. The magnitude of the opposite charges on the two plates must be the same as in the classical picture. Given that the amount of surface charge densities is concluded from the ionic concentration in the segment (in a balanced state), the sum of ionic concentrations must be the same in the segments. A constant electric field is present inside the membrane (across the plates of the condenser), and there is zero electric field inside the parallel conducting plates and outside the condenser. In this ideal picture, the charges are aligned on the border of two (infinitely thin) conducting layers and cannot move. The attractive force between them and the opposite charges on the other plate keeps them fixed in the direction of $z$, and the repulsive force between the charges with the same sign keeps their surface density in the $(x,y)$ plane uniform: the infinitely thin plates are equipotential. This picture illustrates the equilibrium state of charges, consisting of infinitely small non-dissociating charge carriers and perfectly smooth surfaces.
2.4.3 Condenser in dielectric material
In its integral form, Gauss’s flux theorem states that the flux of the electric field out of an arbitrary closed surface is proportional to the electric charge enclosed by the surface, irrespective of how that charge is distributed. The surface layer represents a steep potential and concentration gradient. Above, we assumed that the counterforce that keeps the charges in their place against their electric attraction on the membrane’s surface is a mechanical force: the ions cannot pass through the membrane. However, such a counterforce does not exist on the side toward the ’bulk’ part of the segment. The charges on the plates do not generate an electric field toward the bulk, but the concentration they represent does, according to the Nernst-Planck equation, provided there are charge carriers in the bulk. We hypothesize that virtual electric charges exist in the electrolytes, and their field provides the missing electric field.
When explaining the effect of dielectricity, we must explicitly consider the duality of ions that they obey laws of electricity and thermodynamics simultaneously; furthermore, the complexity of the electric structure of the solution and that the charge carriers have finite size (see "electron size" vs. "dipole size"). According to the theory of electricity, the free charge carriers (the dissociated ions) are located on the surfaces where they form powerfully charged thin layers (condenser’s plates) in the segments separated by a membrane on the two surfaces of the membrane; two proximal charge layers on the surfaces of the membrane. We assume that those ions behave as point charges, i.e., due to the attraction from the opposite charges on the opposite plate, they do not produce an electric field toward the side of the bulk of the electrolyte layers. In the classical picture, those layers represent step-like gradients in the electric field along the $z$ axis, and the constraint that the ions must not enter the membrane provides the needed counterforce.
From a thermodynamical point of view, a driving force acts as long as a concentration gradient between neighboring layers exists. From an electrical point of view, the layers are simple parallel-plate condensers that produce no electric field outside their closed volume and are connected serially; plus, the top layer has free ions. The electric field is proportional to the bulk concentration in the layer. Changes in the local electrical gradient may also alter the degree of dissociation and polarization, producing a graded local electric field. These two driving forces have opposite directions and in a balanced state, the same magnitude. The counterforce acts in a way that constraints the molecules to stay in their layers; it adapts (also compensates for the different interaction speeds) while the gradients are changing. It presses the proximal layer to the membrane and the neighboring layers to each other. (In other words, the internal electric force due to the forcefully changed polarization continuously decreases the electric field and so the potential gradient and so the concentration gradient.)
Due to the membrane’s finite width and surface roughness, the conducting layer has a finite width. The electric field is step-like, with a step size equal to the size of dipoles; furthermore, it is homogeneous in the layers. The rest of that electric field toward the bulk layer directs the dipoles proximal to the neighboring dipole layer, and their polarization increases; in other words, they "produce" an electric field. The final effect is that a dipole layer is attracted to the charged layer, and it forms another layer that shows a charged layer toward the bulk. The process repeats and results in a decaying electric field based on the distance from the membrane’s surface.
In our model, we build the volume of the electrolyte from thin layers of directed dipole molecules (the size of molecules limits the thinness $\Delta z$) in the volume separated into two segments with electrolytes by a membrane with a finite thickness $d$ (that is, a point’s distance in the electrolyte from the two disks will be $z$ and $z+d$, respectively). The contributing potential of an infinitesimal volume on the axis at point $z$ due to the charged sheet is
$dU(c,\Delta z,z)=\underbrace{E^{Gap}_{Classic}(c,\Delta z)*z}_{Potential~{}from~{}sheet}*\frac{1}{z^{2}}=E^{Gap}_{Classic}(c,\Delta z)\frac{1}{z}$
The total potential due to all elements $dz$ from the left side is (in the mathematical formulas below, we express the distance in units of $\frac{z}{d}$, so here $z$ is dimensionless).
[…]
In the case of a single-segment volume, within the segment, a similar potential with an opposite sign is generated by the charges on the neighboring sides of the considered charged sheet, counterbalances the potential described by Eq.(12). However, when a membrane with thickness $d$ separates the segments (with no charges in the gap), the potential on the right side will be
$U_{right}(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl{[}\ln|z+1|\biggr{]}_{1}^{\infty}$ (13)
We can write the electric field in the form
$E(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*ln\biggl{(}\frac{z}{z+1}\biggr{)}$ (14)
That is, the gap sets the potential difference across the membrane to
$U(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl{(}\biggl{[}\ln|z|\biggr{]}_{-\infty}^{0}-\biggl{[}\ln|z+1|\biggr{]}_{1}^{\infty}\biggr{)}$ (15)
We use the approximation that $\ln(\infty)\approx\ln(1+\infty)$, and we arrive at that
$U(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl{[}ln\biggl{(}\frac{z}{z+1}\biggr{)}\biggr{]}_{0}^{1}$ (16)
describes the potential across the plates; that is, the classic potential shall be multiplied by the result of the integral. It was experimentally confirmed [27] that the potential of the membrane is proportional to the logarithm of the distance from the membrane.
2.5 Bridging the micro- and macro views
We assumed that the function can be interpreted for regions $(-\infty,0)$ and $(1,+\infty)$. However, we have arrived at the boundary between quantized and continuous electricity. The minimal thickness of layers is limited, at least by the size of the ion. Furthermore, the membrane’s surface is not flat; there are structures (mainly lipids) with sizes of up to $1~{}nm$, so it is likely realistic to consider a layer thickness of $0.5~{}nm$ as we did above (on our mathematical scale $0.1=\frac{0.5~{}nm}{5~{}nm}$), for which the multiplier is $1.7$. A similar calculation using a layer thickness of $0.1~{}nm$ (on our mathematical scale, 0.02) results in a multiplier of $3.2$; so we use a multiplier of value 3 (that coresponds to a layer thickness of $0.125~{}nm$ (the radius of a $Na^{+}$ ion is $0.116~{}nm$)). Correspondingly, we assume an electric field
$E^{Gap}_{Dielectric}(c,\Delta z)=3*E^{Gap}_{Classic}(c,\Delta z)\quad\biggl{[}\frac{V}{m}\frac{1}{nm*mM}\biggr{]}$ (17)
due to the dielectricity in the segment. This contribution is from the dielectricity in the segment. It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using (pseudocapacitance). By using "roughened anode foil", the thickness of the electrolyte is increased, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
Correspondingly,
$E_{Gap}^{Total}(400,0.5)=4*E^{Gap}_{Classic}(400,0.5)=8.8*10^{6}\bigl{[}\frac{V}{m}\bigr{]}$ (18)
and at a $5\ nm$ membrane thickness, the voltage across the plates becomes
$U_{Gap}^{Total}(400,0.5,5)=E_{Gap}^{Total}(400,0.5)*(5*{10^{-9}})=45\ \bigl{[}mV\bigr{]}$ (19)
"The plasma membrane of all cells, including nerve cells, is approximately 6 to 8 nm thick and consists of a mosaic of lipids and proteins." [1], page 71. Using such a value for the membrane’s thickness may result in a value that is up to 60% higher. Furthermore, the usual concentration is also up to 10% higher. Hodgkin, in 1964, measured molarity values in squid axons for ions $K+$, $Na^{+}$, and $Cl^{-}$, (400,50,40-150) inside and (20,440,560) outside, and they provided potential values $55-75~{}mV$ [25]. Given that we used a plausible but ad-hoc "charged layer thickness" and gap distance, we cannot expect a better agreement. However, our qualitative conclusions remain valid and can be adjusted quantitatively to results of dedicated measurements.
Figure 4: The neuronal membrane’s electric field in the function of the distance from the membrane’s surface. The thickness of the ’atomic layers’ proximal to the membrane’s surfaces are also shown. The electric field is linear inside the uniformly charged atomic ion layer on the surface and is described by Eq.(14) in the electrolyte. For visibility, layer thickness $0.5\ nm$ is assumed. For illustration, a simple capped ion channel in the membrane’s wall is also displayed.
We have the boundary conditions that we know the electric field at the membrane’s surface through the volume charge density, and the Nernst-Planck equation delivers the change of the electric field due to the change in concentration. The physical constraint is that the layer thickness cannot be infinitely small. If we use a parameter for layer thickness equal to that of the surface layer, the calculated electric field distribution will be scalable.
2.6 Ionic currents in a cell
On a microscopic scale, a charge generates a potential field that acts on other charges. It is a common fallacy in physiology that a "potential field travels": neither potential nor current changes without a physical movement of charges. In a conducting wire filled with ions (the solids have a different current transfer mechanism), there are free charges; their number per unit volume is given by $n$, and $q$ is the amount of charge on each carrier. If the conductor has a cross-section of $A$, in the length $dx$ of the wire, we have charge $dQ=q*n*A*dz$. If the charges move with a macroscopic speed $v=\frac{dz}{dt}$, at the macroscopic level, we define the current $I$ as the charge moved per unit of time as
$I=\frac{dQ}{dt}=q*n*A*v$ (20)
The effect was hypothesized at the very beginning: "it seems difficult to escape the conclusion that the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment." "it is necessary to suppose that there are more carriers and that they react or move more slowly" [28]. The charged ions, attracted by the opposite charges on the other plate and repelled toward the only exit point AIS really moves slowly and behaves as if it were a component of the membrane.
Notice that if any of the factors is zero, the macroscopic current $I$ is zero. Microscopic carriers must be present in the volume, and have charge, the cross section must not be zero, and the charge carriers must move with a potential-assisted speed, which needs an electric field generated by the system, see Eq.(7). When there is no constraint force, the thermodynamic and electric forces will result in an effective force moving the ions. That is, the ions can move without an external electric field, as experienced in biological systems. One of the fundamental mistakes by HH was neglecting the Coulomb force for ion’s electric interaction and that the electrical potential is created by diffusion processes instead of ideal electric batteries.
In biological systems, where ions represent the "current", it is either a native current (without an external potential) or an artificial injected current or potential generating a current. This way, the current consistently produces or is accompanied by a change in concentration gradient, as the moving ions represent both mass transfer and charge transfer simultaneously. The potential and current are connected through the features of the medium (material) that hosts our measurement.
According to Stokes’ Law, to move a spherical object with radius $a$ in a fluid having dynamic viscosity $\eta$, we need a force
$F_{d}=6*\pi*\eta*a*v$ (21)
(drag force) acting on it. A (microscopic) electric force field $\frac{dV}{dz}$ inside the wire would accelerate the charge carriers continuously
$F_{electric}=k*E(z)*q$ (22)
with a constant speed $v$. It is not the drift speed because of the electric repulsion, it is a potential-assisted or potential-accelerated speed that can be by orders of magnitude higher. The medium in which the charge moves shows a (macroscopic, speed-dependent) counterforce $F_{d}$, which in steady state equals $F_{electric}$, that is :
$I=\frac{k*q^{2}*A}{6*\pi*\eta*a}*n*E(z)$ (23)
We extend the original idea by using
$E(z)=E_{electric}+E_{thermal}+E_{xternal}\quad in\ electrolytes$
The amount of current in a wire is not only influenced by the electric force field (specific resistance) but also by the number of charge carriers $n$. While the latter is commonly considered constant and part of the former, this is not necessarily the case for biological systems with electrically active structures inside. The medium’s internal structure introduces significant modifications (according to E. Schrödinger: "because the construction is different from anything we have yet tested in the physical laboratory" [20]). Applying an additional electric field to a wire can generate varying amounts of current as the number of charge carriers changes due to biological effects. For axons, we use a single-degree-of-freedom system, a viscous damping model, so the ions will move with a field-dependent constant velocity in the electric space; so it takes time while they appear on the membrane. The activity of potential-controlled ion channels in its wall may change $n$ in various ways; furthermore, that change can result in ’delayed’ currents during measurement, for example, in clamping, as physiological measurements witness.
If we have a concentration $C(z)$, in the volume $A*dz$, we have $dQ=C(z)*A*dx*q$ charge, resulting in another expression for the current
$I=\frac{C_{k}(z)*A*dz*q}{dt}=C_{k}(z)*A*q*v$ (24)
By combining equations (23) and (24):
$v(z)=\frac{k*q}{C_{k}(z)*\eta*6*\pi*a}*E(z)$ (25)
The higher the resulting space derivative (gradient) and the fewer ions that can share the task of providing a current, the higher the speed. We hypothesized (it needs a detailed simulation) that in the case of this charged fluid, the electric repulsion plays the role of ’viscosity’ [7]. The higher the charge density, the stronger the force equalizing the potential; so $\eta$ is the lower, the higher the charge density (proportional to $C_{k}$). For the sake of simplicity, we assume that the speed is proportional to the space gradient of the local concentration and potential. Recall that our equations refer to local gradients only. The electric gradient can propagate only with the speed of the concentration gradient, given that only the chemically moved ions can mediate the electric field [7]. The lower interaction speed limits the other interaction speed if the interactions generate each other.
The dependence of the diffusion coefficient on the viscosity can be modeled by the Stokes-Einstein relation:
$D=\frac{k*T}{6*\pi*\eta*a}$ (26)
so we can express the speed with the diffusion coefficient $D$ and temperature $T$
$v(z)=\frac{D}{T}*\frac{q}{C_{k}(z)}*E(z)$ (27)
Models in neuroscience (as reviewed in [29]) almost entirely ignore these aspects. In our physical model, we see that the measurable membrane potential and current change in the function of the ions’ speed, the concentration, and its time derivative; furthermore, all mentioned quantities depend on the effective potential.
3 Forming membrane’s potential
We considered that the neuron has a stable base state. On the one side, this resting state must be dynamically stabilized for little perturbations (and to provide a mechanism when the cell grows, divides, or ages). On the other side, it must be able to restore the state after rough perturbations, causing short-time transients (when restoring the membrane’s potential after issuing a spike). The two states need different physical mechanisms.
3.1 Neuron as system controller
From control theory, it is known that the goal of the system is to govern the application of system inputs to drive the system to a desired state while minimizing any delay, overshoot, or steady-state error and ensuring control stability. The neuron implements a controller that monitors the controlled process variable (membrane voltage) and compares it with the reference or setpoint (the resting potential). The difference between the process variable’s actual and desired values called the error signal, is the actual offset potential. It is applied as feedback to generate a control action to bring the controlled process variable to the same value as the set point.
Biology uses such a simple controller. The gradients are used to adjust the process variable by their positive and negative contributions, and the different speeds of the thermodynamical and electrical interactions minimize the delay (i.e., provide the maximum speed of operations, vital for survival). The steady-state error is minimized by setting the process variable to the reference point using long-term stable parameters (geometry and overall concentration). The low-intensity current through the always-open resting ion channels provides dynamic stability in the steady state. The permanently zero error signal indicates a lack of operations. In contrast, the permanently nonzero error signal is a sign of abnormal operation and is likely a symptom of a neurological disease.
The system aims to provide an intense output that informs the downstream neurons. This overshoot is implemented by a large number of voltage-gated ion channels (not identical with the non-gated ion channels used to maintain the resting potential) distributed in the neuron’s wall. The overshoot current flows through a similarly large number of persistently open ion channels concentrated in AIS (it has about two orders of magnitude higher density of channels than that in the wall). The intense slow current produces a condenser-like behavior (capacitive current); the phenomena called polarization and hyperpolarization (not polarization; instead, a movement of completely separated charge in the surface layer of the electrolyte) of the membrane provide the necessary positive and negative error signals for the controller in the transient state by moving the actual potential value of the membrane above and below the resting potential value.
From a control theoretical point of view, the condenser geometry sets the ’always the same’ value (the set-point) and the currents serve to keep or restore that value. Forming membrane’s potential is an excellent example of the principle that nature uses "one process to build and the other to correct" [6].
3.1.1 Biology builds
After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq.(19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, we discuss only the ’downhill’ channels), the significant difference in current intensity needs a different discussion.
3.1.2 Electricity and thermodynamics correct
The potential’s and the electric field’s magnitudes depend on the concentration and the geometry (the finite width of the membrane). The ions’ chemical nature comes into play only if the polarizability can differ for different molecules.
Given that the same number of charged ions must be present on the two sides of the membrane, their resultant surface density $\sigma_{A}$ must be the same. Notice that the effect is purely electrostatic, resulting in an asymmetric ion distribution; no permeability is needed. If the membrane is (slightly) permeable, ions will move across the membrane until equilibrium is reached. The resulting potential difference is directly proportional to the concentration difference.
The classical condenser has an electric field, as the blue dashed diagram line shows in Fig. 4. No field is outside or inside the condenser plates (see Fig. 3). There is a sudden jump in the field on the conductor’s surface (the armatures) and a constant field between the plates. (The charged layer can be ideally thin if electrons are sitting on the surface but is of finite thickness on the opposite side where ions are on the surface. This latter effect is not discussed in classical electricity.)
The biological condenser behaves differently; see the red diagram line. The electrolyte (instead of an insulator) outside the condenser drastically changes the structure of the field. The electric field on the surface and between the plates changes by a factor of 3 to 4. Outside the plates, the polarization creates a field that changes logarithmically. An ion layer with finite thickness is built near the membrane’s surface. It is uniformly charged, so the field inside it is linear. At the internal surface of the condenser, it takes the value of the field calculated for the gap; on the other side of the layer, it takes the value of the logarithmic field at that position. The dashed line represents the gap field, the continuous line represents the charged layer field, and the dotted line represents the electrolyte field. We can observe the "correspondence principle": the particle-continuous fields join seamlessly; furthermore, if the polarizability of the electrolyte decreases toward zero, the field value approaches the classical value.
The thickness of the charged layer has a significant impact on the resulting electric field (i.e., the neuron’s ’setpoint’). If biological fragments are formed and settle down on the membrane’s intracellular side, due to the increased $\Delta z$, the resting potential (and the threshold potential) may change, leading to neurological diseases.
3.2 States
The two electric fields show essentially different dependencies on the neuron’s parameters. Recall that although the parameters we used are biologically plausible, they are only estimations. The qualitative conclusions, however, remain valid. As Figure 1 shows, in the function of the membrane’s thickness, the electric field across the membrane due to electric charges (it does not depend on the thickness of the membrane), for different physically plausible assumptions; furthermore, the thermal electric field.
The balanced state is set to where the thermodynamic and electric "electric field" diagram lines cross each other; around $5\ nm$; the value we assumed in evaluating our equations. On the other hand, we can check the electric field’s dependence on the summed-up concentration of the ions in the segment (see Fig. 5). Here, the thermodynamic electric field is constant but different at different width parameters. The classic condenser does not provide a reasonably balanced state (a matching line). The dielectric diagram line seems to provide a realistic estimation when using biologically reasonable parameters. One can describe the neuron as a balanced system at the intersection of the sloping and the horizontal lines. In the resting state, only minor deviations from this state exist in both the electric field and the chemical concentration.
Figure 5: The electric and the thermal "electric fields" at different concentrations, see Eq.(18).The thermal electric field varies with the thickness parameter and depends on the concentration; see Eq.(4). Balanced states are where the sloping lines cross the horizontal ones. The dot marks where HH provided measured data.
In the transient state, if we assume that a charge transfer causes a significant potential change around $100\ mV$ in the layer of width $100\ nm$, it means a $0.1*10^{6}\bigl{[}\frac{V}{m}\bigr{]}$ jump in the electric field that corresponds to a $100\bigl{[}mM\bigr{]}$ jump in the concentration (for only a short time and only in that mentioned layer not in the bulk; see also Fig. 7). During issuing an AP, the system finds its way back to the crossing point (the path is not simple: both concentrations on both sides of the membrane and the electric field in the mentioned layer change; the process is relaxation, as it is seen in the shape of an AP).
Even we can imagine how the electric process happens in this controlled system (for visibility, the positions of the marks are not proportional to the mentioned numbers). The system is in a balanced state at a crossing point. The two circles connected by bent arrows represent that when an excitation begins at the point with lower concentration and higher potential (the cytoplasm side and negative resting potential), ions rush into the surface layer, so the ion concentration increases, that increases the electric field; which is seen that the membrane potential suddenly increases. The increased potential starts a current, but the current is slow, and the intensity of the current through the channels in the membrane’s wall is higher than the current of the drain (AIS) so the system issues an AP while returning to its starting point. Given that it receives a large amount of $Na^{+}$ ions, its primary effort directs to get rid off them (the electric gradient cannot distinguish the electric charges by their chemical nature), so the output current through the AIS fundamentally comprises $Na^{+}$ ions. The system will move along the line, with the restriction that it is a movement in the phase space (the change in concentration is not proportional with the time), and the different points on the membrane’s surface may have the potential at different times (the spatiotemporal behavior). Due to the capacitive current, the local potential may temporarily drop below the set-point (as observed in physiology, the phenomenon is known as hyperpolarization). The time course of the processes are described by the laws of motion of biology [7] (the time derivatives of the Nernst-Planck equation), and the simplified model works as described in section 4 of [30] and its computational details in [31].
3.2.1 Resting state
In the resting state, the goal is to provide stability by using slow and less intense currents. As [1] discusses, there are ’holes’ (non-controlled ion channels) in the membrane where the counterforce is missing. The difference in the electric and thermodynamic forces could accelerate ions until the Stokes-Einstein speed (see Eq.(27)) is reached. Given that the deviation from the set-point is slight, the intensity and speed of the current are low. The slow ion transport delivers both charge and mass, so the gradients permanently direct the process control variable toward the reference point. Given that the ’resting channels’ are scattered on the surface of the membrane and it takes time for the gradients delivered by slow ions to reach the entrance of an ion channel (see also section 4), some slight variations in the actual value of the membrane’s potential necessarily exist. Given that the AIS comprises non-gated ion channels with much higher density than those in the membrane, the variations and the control ("leaking") are marginal, and for mainly through the ion channels in the membrane’s walls, as observed in experimental physiology.
The electric voltage contribution is unidirectional, the two-component thermodynamic contributions, depending on the concentrations, can be positive or negative. If choosing the zero potential at the middle of the electric potential, we can interpret that the concentration combinations maintain their ionic concentrations (and so also thermodynamic voltage contributions) independently on both sides starting from a zero potential level. They generate potentials autonomously in the segments proximal to the opposite surfaces of the membrane. The electric charge on the membrane creates another voltage independently. However, that voltage must fit between the two thermodynamic voltages. Any deviation in the potentials starts a current directing the concentrations and voltages toward a balanced state. The process is complex: the sum of surface concentration of ions must be the same on both sides for the condenser and the ratios on the intracellular and extracellular sides must be adjusted to satisfy the conditions
$U_{internal}^{K^{+},Na^{+}}$ $=-U_{electric}$ $+U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}}$ $(+U_{internal}^{invasion})$ (28)
$U_{external}^{Cl^{-},Ca^{2+}}$ $=U_{electric}$ $-U_{external}^{Cl^{-}}$ $-U_{external}^{Ca^{2+}}$ $(+U_{external}^{invasion})$ (29)
In a balanced state, the left sides are zero: at the contact point between the electrolyte and the membrane there are no potential differences, while in perturbed state they are non-zero and so they represent the driving forces for restoring the balance. (Recall that the ions are slow and that the local electric and concentration gradients control the local potentials.) From the derivation, follows that
$U_{electric}$ $+\displaystyle=-2*(U_{internal}^{K^{+}}$$+U_{internal}^{Na^{+}})$ (30)
$U_{electric}$ $=-2*(U_{external}^{Cl^{-}}$ $+U_{external}^{Ca^{2+}})$ (31)
$U_{electric}$ $=U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}}$ $=-(U_{external}^{Cl^{-}}$ $+U_{external}^{Ca^{2+}})\quad$ (32)
Figure 6: The mechanism for controlling neuronal operation is the same in all cases shown in Table 1. It appears that the thicker membrane produces a higher membrane electric potential and requires a higher thermal potential to compensate for it. However, it can solve the task using a lower salt concentration. Given that the electric membrane potential accelerates the ions across the ion channels in the membrane, and the operation is faster. The electric field has a similar value, as shown in Fig. 1. Furthermore, the total concentration decreases as the membrane thickness and potential increase.
The membrane’s electric potential defines the resting potential, and the "electric" and "thermodynamic" potentials work together to control neuronal operation. The anatomically defined "electric" membrane potential plus the (compared to the internal amount) vast amount of ions in the extracellular segment and so the stable concentration provides a solid reference point. In the resting state, the membrane potential is balanced the presence of always-open "resting ion channels". When ions flow into the membrane, the "resting ion channels" can keep the balance: the internal $K^{+}$ and $Na^{+}$ concentrations are adjusted as the $Na^{+}$ flows in.
Figure 6 also shows that in its native state, the neuron is electrically balanced: in the internal and external segments, the resulting thermodynamical and electrical forces are equal. Any foreign invasion into the neuron’s life (chemical or electrical, including clamping) changes the left-side reference value to an externally set value. The neuron must adjust its $Na^{+}$ and $K^{+}$ concentrations to reach a new balance while the invasion persists. That is, any invasion ("testing in the physical laboratory") changes the concentrations and voltages inside the cell. In addition, the internal biological processes can trigger internal processes that contribute (from the measurement’s point of view) foreign currents or processes, see also section 3.2.2. The feedback the clamping methods apply introduces a compensating current. Even a simple conductance measurement device is active: it applies a test voltage that generates a test current, which sensitively affects the ionic composition of the neuron (recall that the concentrations may differ by up to six orders os magnitude). Although modern electronic devices attempt to conceal this effect, they subtly influence biological operation.
From our result, it follows that for more chemical elements, a per-element coupled set of equations exists plus the sum concentration agreed electric fields that are valid simultaneously and define the process variable (aka membrane potential) of the neural regulatory circuit. In a more symmetrical form
00 0 $=\frac{U_{electric}^{Total}(c,\Delta z,d)}{2}+\sum_{k}U_{thermal}^{C_{k}}(d)+{invasion}\quad for\ all\ k;\ inside\ and\ outside$ (33)
$c$ $=\sum_{k}C_{k}^{ext}=\sum_{k}C_{k}^{int}$ (34)
That is, we have an unidirectional electric potential difference and oppositely directed thermal voltages. In the balanced state, the two potentials must be equal. When changing say one concentration, it changes the other concentration on the same side, and the surface charge density of the condenser. The changed electric voltage re-adjusts the thermodynamic voltage, and as a consequence, the concentrations, in the other segment. There is no simple way to express, say, $[Na^{+}]_{i}$ in the function of $[K^{+}]_{i}$ or vice versa, and similarly for the other ions. Everything moves. The size matters: the amount of ions in the external segment is many orders of magnitude higher, and the gradients are reciprocally proportional to the amount of ions.
As the figure suggests, an external invasion, typically, an electric voltage on the intracellular side, changes the balanced state and due to the parameters are linked, changes all concentrations. When moving the system out of its balanced state, in any way, a driving force appears that moves the system towards finding a new balanced state. However, recall that the ions are slow, so the changes are not instant.
Table 1 shows our results for three different famous published cases. It looks like the calibration of the concentration of negative and positive ions needs scaling and the measurement of $Ca^{2+}$ comprises issues.
Table 1: The summary data table for equilibrium (Data from Table 2.1 in[25] )
Ion Inside Outside $U_{therm}\frac{RT}{zF}log\bigl{(}\frac{[C]_{o}}{[C]_{i}}\bigr{)}$ $U_{electr}^{theor}$ $U_{electr}^{exper}$ $U_{membr}$ Width $E_{membr}$
[mM] [mM] [mV] [mV] [mV] [mV] [nm] [MV/m]
Squid (HH)  T=293 $K^{o}$ 3 14.0
$K^{+}$ 400 20 $-75.5=58log\bigl{(}\frac{20}{400}\bigr{)}$ 41.9 41.4 -62.0
$Na^{+}$ 50 440 $+54.8=58log\bigl{(}\frac{440}{50}\bigr{)}$
$Cl^{-}$ 40 560 $-66.5=58log\bigl{(}\frac{560}{40}\bigr{)}$ 41.9 42.0 63.0
$Ca^{2+}$ $0.27^{*}$ 10 $+45.5=29log\bigl{(}\frac{10}{0.27}\bigr{)}$
used for the published value 0.4
Frog muscle (Conway)  T=293 $K^{o}$ 8 9.61
$K^{+}$ 124 2.25 $-101=58log\bigl{(}\frac{2.25}{124}\bigr{)}$ 88.9 83.6 -125.4
$Na^{+}$ 10.4 109 $+59.2=58log\bigl{(}\frac{109}{10.4}\bigr{)}$
$Cl^{-}$ 1.5 77.5 $-99.4=-58log\bigl{(}\frac{77.5}{1.5}\bigr{)}$ 88.9 82.7 124.1
$Ca^{2+}$ $0.021^{*}$ 2.1 $+58.0=29log\bigl{(}\frac{2.5}{0.25}\bigr{)}$
used for the published value 4.9
Typical mammalian cell  T=310 $K^{o}$ 6 11.1
$K^{+}$ 140 5 $-89.7=62log\bigl{(}\frac{5}{140}\bigr{)}$ 57.7 57.3 -85.8
$Na^{+}$ 15 145 $+61.1=62log\bigl{(}\frac{145}{15}\bigr{)}$
$Cl^{-}$ 4 110 $-89.2=-62log\bigl{(}\frac{110}{4}\bigr{)}$ 57.7 56.9 85.9
$Ca^{2+}$ $0.04^{*}$ 5 $+60.8=31log\bigl{(}\frac{5}{0.04}\bigr{)}$
used for the published value 0.0001
$Na^{+}$ ions only [4] T=310 $K^{o}$ 6 10.0
$Na^{+}$ 12 145 $+66.6=62log\bigl{(}\frac{145}{12}\bigr{)}$ 60.7 66.6
When calculating the values in Table 1, we used $\Delta z=1.05$ and employed the membrane’s thickness data from various publications to demonstrate that our theoretical approach is correct and that using the correct data can provide even the absolute values of the membrane potential. The higher-than-expected value of $\Delta z$ may suggest that the ions and membrane’s double layers in the electrolyte may also play a role. Dedicated, complete investigations can reveal the details.
3.2.2 Transient state
Figure 7: The mechanism of producing AP (using numbers referring to the case of squid shown in Table 1). In the resting state, the inside thermal potentials follow the green paths. When the rush-in of $Na+$ ions suddenly increases the internal concentration to $[200]$, the potential on the internal side of the membrane increases to $+90\ mV$, and the neuron must decrease its $K^{+}$ concentration to restore the resting potential, or release an action potential through the AIS.
In a special type of invasion, when suddenly a large amount of $Na^{+}$ rushes in (i.e., the trigger is external, but the ions derive from a biological process), the current throughput of the "resting ion channels" is not sufficient anymore: the internal $Na^{+}$ concentration drastically increases, and so does the membrane’s potential. At that point, the neuron is out of balance: although most of the $Na^{+}$ ions can flow out through the AIS, the system also changes its $K^{+}$ concentration through the limited capacity "resting ion channels" until the balanced state regained. This latter effect must not be mismatched with the capacitive current, which (due to its changed direction) causes the phenomenon known as "hyperpolarization".
Compare the figure to Fig. 7.2 in [1]. Notice the two-sided negative feedback effect: a change in the value of the "command potential" or "command concentration" changes the other value. In the case of voltage invasion, the ratio of the ion concentrations also changes. In this case, the feedback amplifier is represented by the delicate balance of the electric and thermodynamic force fields. In the case of clamping, another (electric) feedback amplifier is introduced into the system and they work against each other. The neuron attempts to restore its balance without external potential (it cannot distinguish the foreign potential from its membrane potential), and changes its concentrations accordingly. A vital difference is the speed of the feedback amplifier. The electrical one works with electric speed, while the thermodynamic one with the several orders lower thermodynamic speed. When switching (or changing) the external ("command") voltage, the neuron must first get into its "steady state". It takes time. "The command potential, which is selected by the experimenter and can be of any desired amplitude and waveform" [1]. However, if the measured data readings are not from a steady state, the measurement is wrong: the different amplifier speeds (different current speeds) surely distort the measured value.
Figure 7 shows the schematic course of potentials and concentrations during issuing an AP. Follow the green path: the concentration ratio of $Na^{+}$ and $K^{+}$ ions define a thermodynamic contribution in the intracellular segment. Follow the blue path: the $Cl^{-}$ and $Ca^{2+}$ ions define a thermodynamic contribution in the extracellular segment.
On the left side, two positive ions are in the solution, and the membrane potential (due to charge separation) rises toward the outside segment; the ions cannot go "uphill". Similarly, the $Cl^{-}$ ions cannot go uphill (due to their negative charge, the same potential is also rising for them). This single membrane potential keeps all ions in their segments. However, the $Ca^{2+}$ ions can travel "downhill". Calcium pumps needed: although in minimal concentration, $Ca^{2+}$ ions flow into the intracellular segment, and they must be pumped out. (BTW: this is likely also the reason why measuring $Ca^{2+}$ concentration is inaccurate.)
As we discussed above, the concentrations and voltages in the extracellular segment, with its vast amount of ions, remain unchanged in resting and transient states. Follow the red path: when $Na^{+}$ ions rush into the intracellular segment; they increase the concentration from $[50\ mM]$ to (say) $[200\ mM]$ (from point (Rest) to point (AP)). This sudden change increases the thermal contribution of $Na^{+}$ potential to $[-20\ mV]$, but because the total charge on the membrane significantly increases, the resulting voltage on the intracellular side of the membrane increases to $[90\ mV]$. The neuron could decrease its potential to the resting value if it could decrease $K^{+}$ concentration to $[100\ mM]$. However, the ion delivering capacity of the "resting ion channels" is not sufficient: they are sized to maintain the resting state. (Furthermore, the electric contribution appears instantly, while the thermodynamic one, due to the finite speed of ions, only with some delay [7].) Follow the dashed orange path: the large amount of rush-in $Na^{+}$ ions drastically increases the surface concentration on the membrane, so the potential suddenly changes to $[90\ mV]$. Follow the dotted orange path: the electric field across the membrane suddenly changes its direction from positive to negative (see the slopes of the solid and dotted orange lines) so that the excess positive ions can move "downhill" out of the intracellular space through the high-capacity ion channel array (AIS) at the end of the neuron. Its internal end is at the high membrane potential, while the external end is at the low potential of the extracellular segment.
The set-point on the extracellular side remains fixed. On the intracellular side, at the beginning of the AP, the actual voltage is above that of the set-point on the extracellular side, so at the two ends of the AIS, a positive driving force moves the ions toward the downstream neuron. Given that positive $Na^{+}$ ions leave the intracellular segment, its potential continuously decreases. At some point, the voltage drops below the set-point of the extracellular side and the electric field across the membrane reverses. The driving force also decreases, and a low-intensity current restores the balanced state, to the previous set-point. When measuring at the AIS, one observes that the potential rises suddenly, then decreases below the set-point (hyperpolarizes the membrane), then a decreasing current (flowing in the opposite direction due to the changed slope of the resulting potential (the changed electric field across the membrane), but comprising the originally rushed-in $Na^{+}$ ions) slowly restores the resting potential. Here comes to light the finite speed of ions again. The ions can follow the potential changes with a delay due to their finite speed, causing an apparent delay in the time course of the AP’s current. Until they travel, the field at the AIS changes, but the finite speed of the ions cannot deliver that field in time. The current "overshoots" then relaxes. This is the normal operation of an $RC$ circuit; the capacitive current is misinterpreted as delayed $K^{+}$ current.
3.3 Mathematics of the transient process
During regular operation, a very intense current pulse is created: "A small flow of ions carries sufficient charge to cause a large change in the membrane potential. The ions that give rise to the membrane potential lie in a thin ($<1\ nm$) surface layer close to the membrane" [26] (see the dotted line in Fig. 4; notice that the horizontal scale on the figure spans a few $nm$). For this, a considerable number of ion channels are used. When the trigger signal is received, the gated ion channels in the membrane’s wall open quickly and only for a brief period. The rushed-in ions create a large offset potential in the layer proximate to the membrane’s surface (from an electrical point of view, a condenser) that drives an intense current through the always-open ion channel array AIS (from an electrical point of view, a resistor) at the beginning of the axon. The slow current on the membrane’s surface mimics a condenser: in the picture of classical electricity, the charges stay inside the distributed condenser while they travel on its surface. In classical electronics, the condenser is sizeless.
Those two elements form a serial $RC$ circuit that acts as a damped oscillator and produces a single current pulse, known as AP. Neuroscience mistakenly claims it is a parallel $RC$ circuit. As discussed in the previous subsection, there are ’resting’ ion channels in the membrane’s wall, and they are in parallel with the resistor. However, their number and intensity are insignificant for producing AP. In this action, the gated ion channels in the membrane’s wall produce an instant pulse that appears as a peaked current time course due to the slowness of the charge carriers. The output current flows as long as the offset potential persists.
The process can be modeled as follows: a sudden voltage gradient appears on the membrane (as a discrete capacitor), and a current flows out through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [28] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see [30], section 4; furthermore, for further details [14]. As we discussed there, the capacitive current, which per definitionem changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have enough transmission capacity, although some $K^{+}$ current surely exists.
The shape of the output waveform depends on the pulse width ratio to the $RC$ time constant. When $RC$ is much larger than the pulse width, the output waveform resembles the input signal, even with a square wave input. (In the case of the neuronal oscillator, the time constant is about $\zeta=0.35$, so the shape of the front side of the spike is similar to the one derived for the rush-in current, while the back side is very prolonged.)
3.4 Robustness of the control circle
An exciting thought experiment is discussed in [4]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $[Na^{+}]_{o}$, the outside or extracellular sodium concentration, is $145\ mM$, and $[Na^{+}]_{i}$ is 12 mM. And suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential sets up to $66.6\ mV$. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In lack of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law, alone, does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates its electric potential, which is a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electric potential. However, the thermal driving force exists, and the started ion transport quickly builds the electric layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage remains the same, and the voltage defined in this way adjusts the concentrations appropriately. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form a membrane, establishing cellular electricity. As Fig. 6 depicts, the concentrations may adapt to the living conditions (whether living in see-water), but the general operating principles remained the same. The see-water environment enables using thinner membrane and lower membrane voltage; however, the same electric field is is used in all cases.
Given that neurons share the external (global) concentrations (defined by the vast amount of ions in bulk), they provide a firmly fixed offset value to the always-the-same electric potential. This way, the operations of the individual neurons do not interfere. Furthermore, the internal (local) concentrations can adjust themselves even following a very rough charge perturbation while issuing an AP; moreover, when the living organism is growing.
3.5 Leakage current
Although through $\Delta z$, the ion’s water-related behavior can influence the surface charge density, one can draw the conclusion that when only one type of ion is solved in the liquid, approximately the same concentration difference between the segments produces approximately the same "thermodynamic electric field" that can counterbalance the electric field of the biological condenser. One can notice that in the classic measurement by Hodgkin-Huxley, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) are 380 and 390, respectively, while the sum concentrations of the $Cl^{-}$ and $Ca^{2+}$ are 450 and 460; practically all equal, as our analysis derived it must be. The deviations are well within the uncertainty of the measured values.
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage is generated. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a $42.5\ mV$ voltage [3]. In HH’s picture, the question what ion constitutes the current remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leak^{+/-}$ ions (??) (depending on the direction, it could be positive or negative), and neither of them changes the concentration neither on the departure nor on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law.
The idea of leakage current was wrong: the voltage HH measured [3] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on its sides, even when no current flows. Measuring the energy consumption of neuronal operation [13] confirms that no resting current, in the sense as HH used the notion, exists. The existence of leakage current would be against the overall metabolic efficiency of neuronal operation.
3.6 Heat emission/absorption
The lack of resting current also solves the long-standing mystery of "reversible heat release during the action potential", which is inconsistent with the Hodgkin–Huxley model. In the framework of that theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}*R$) the ion currents flow. As Hodgkin wrote [32]: "Hill and his colleagues found [33] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation."
The issue here is that the dissipation was attributed to the "leakage current." (Although, due to the different conduction mechanisms, an ion current is not dissipated in the way electron energy is dissipated in solids.) However, as our explanation shows, the voltage generated across the membrane is due to charging rather than a permanently flowing current, so no dissipation was present given that no resting current was present. Although the energy to build a potential on the membrane and to maintain it needs biological energy consumption, the process of generating an action potential is, in principle, reversible: ions flow into the layer on the membrane and flow out through the AIS.
According to that, the "temperature of individual particles" with the mean kinetic energy $E$ is $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is the temperature is directly proportional to the kinetic energy. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electric potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature. In the second phase of the AP, those more energetic ions that provide excess voltage above the resting potential leave the membrane through the AIS, so the temperature decreases in the second phase. The heat liberation and absorption are a direct experimental evidence of the "slow current" and the mechanisms described above.
3.7 Goldman-Hodgkin-Katz potential
In our clear physical picture, the thermodynamical forces on one side of the membrane and on the other side are summed and balance each other. As we emphasized, the concentration gradients are ion specific; furthermore, that the Nernst equation comprises a per-ion indefinite constant. The absolute concentration alone has no meaning for a thermodynamical gradient; furthermore, one cannot find a physical reason why concentration on one side should be intermixed with concentrations on the other side. Combining non-gradient type meaningless concentrations in an expression used for deriving electric gradient is not more than number magic; it has no physical background. Even if we intermix arbitrarily chosen ions from both sides and multiply by a logarithm base. The idea itself is a nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability nor with ion absorption [18]. The causality is reversed: the potential is a static notion and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of forming the gradients, mobility and permeability play a role but not in determining the concentrations or the resting potential.
The $Ca^{2+}$ ions do not fit into the GHK picture. However (see our Table 1), it provides a thermodynamic contribution in the same order of magnitude as the charge separation by the membrane. If we add the potential due to $Ca^{2+}$, we arrive at the correct result that the total balance of the thermodynamic contributions equals zero, that is, the resting potential is zero. The electric potential contribution is an additional term, resulting in the resting potential.
It was already stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions adsorption on the membrane surface generates the membrane potential", for a review see [2]. In the light of our findings, we must reconsider a few key concepts. We derived (see Eq.(19)) that the gap voltage depends on the thickness of the membrane and the overall concentration. Furthermore, we demonstrated that the gap voltage is equivalent to the resting potential (correctly: potential difference) and is independent of ion mobility or ion composition in the solution.
When adding a new ion to the solution, the sum concentration increases, and so the electric force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$). No linear summing or similar assumptions can be made. The potential is described by coupled equations as discussed in connection with Eq.(34). As we discussed in connection with Eq.(2), the integral form of the Nernst equation implies an undetermined additive constant. To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is changing the base of the logarithm used in a calculation to match the experimental value.
3.8 Reversal potential
From our picture, it is clear that the reversal potential is reached when the potential on the inner segment is lowered or lifted in such a way that the orange sloping potential in Fig. 7 turns into horizontal. To reach it (provided that the potential on the other side is kept fixed), the concentrations on the invasion site must be changed. The invasions resulting in the exact (but opposite) change of potentials from the two sides leads to different concentrations. The invasion acts as changing the potential at the corresponding surface side of the membrane. Given that the difference between the two potentials on the two surfaces defines the driving forces, the concentration on the other side is also affected. That is, the reversal potential exists, but its value depends on the context, see the discussion about invasions.
4 Moving in the membrane’s potential
In Fig. 4, an ion channel is depicted in the middle of the figure with a diameter of about $3\ nm$ for visibility. Furthermore, we assume that the ion’s size and, correspondingly, the thickness of the atomic layer in the electrolyte on the surface of the membrane is about $0.1\ nm$ (although with their electric double layers, they can also reach the size of $1\ nm$). For comparison, recall that the size of the tip of the clamp pipette is in the range of $1,000\ nm$ and the size of the soma in the range of $10,000\ nm$.
4.1 Passing through ion channels
Figure 4 also hints at how ions can move in the proximity of the membrane. On the left side, the electric field increases toward the membrane, so the ions cannot move in that direction without an initial force being applied. The case is the same on the right side, because the charge is the opposite. An ion must gain energy to get closer to the membrane, which means, at the same time, higher potential energy. This situation explains why the ions do not diffuse into the bulk region. At their present positions, the counterforces by the membrane (and the dipole layer on it) balance the attraction from the opposite side.
Notice that the resulting electric field cannot accelerate an ion except when the ion is at the beginning of the open ion channel across the membrane. Here, the mechanical counterforce is missing, and the field drops to the value generated by the condenser inside the plates. Here, a vast electric field accelerates the ions to a high speed (although a force needed to move against the viscosity decreases the electrostatic acceleration). The empty red circles show an ion that traverses through the channel. The figure shows their electric fields on the two sides, at the beginning and end of the channel. (On the right side, the electric field is negative, but so is the charge of the ions.)
4.2 Moving in the polarized electrolyte
The electric acceleration sharply decreases after leaving the ion channel (see the right side). As shown by the difference of the vertical positions of the filled red and green balls, the difference in the electric field values is huge across the membrane ($E_{accelerate}$) and considerable inside the dielectric electrolyte ($E_{assist}$) as shown by the difference in the electric field values of the green balls. The field still accelerates, but significantly less intensely than the one in the gap. Suppose we assume that the ion quickly accelerates to its Stokes-Einstein speed. In that case, the travel speeds in those space regions are proportional to the difference of the fields in the shown positions and the reciprocals of their travel times.
The passed ions decrease the concentration and the electric field on the high-concentration side, and, concersely, increase them on the low-concentration side. Similarly, the local gradients decrease and increase. At the exit, the field continues to accelerate the ion. However, the increasing potential and concentration increasingly decelerate the passing ions, so they quickly brake them to the assisted speed. Then, the effect of the electric field cancels, and the ions will move with their diffusion speed ("the ion stops"). The local potentials, connected to the operation of the ion channels, act as an "electric lock".
4.3 Ion selectivity
As Eq.(6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamical forces are equal. The $E_{thermal}^{C_{k}}(d)$electric field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions are different. While one ion experiences a resulting force and moves to the other segment, the other remains resting. Of course, the ions moving across the segments change the overall concentrations in both segments (and so the electric field contribution), indirectly affecting the balanced state of the other ions. When a mixture of ions is present in the volume, all ions feel a different driving force, and the solution as a whole will be balanced when the resulting forces for all chemical ions are balanced. The permanently open ion channels only passively participate in the process. The driving forces of the ions are different, and they produce, depending on the actual concentrations, the selectivity of the channels.
When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential. All other ions, including $K^{+}$, feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave controversially during the transient period. The actual voltage gradient may temporarily direct the chemical gradient in opposite directions. Phenomena such as ’over-relaxation’, in this sense, similar to ’hyperpolarization’, may happen.
However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or a few atomic layers on top of each other, depending on the concentration and voltage in the bulk on the two sides. (The voltage gradient diagram line validates the plane crossing the membrane and the ion channel.)
Our results align with the observation (see caption of 11.22 in [26]). The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a large change in the membrane potential. Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions.
5 Summary
By introducing theoretically that biological currents are slow we underpinned the experimental observation by Hodgkin and Huxley [34] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Recognizing that biological systems use charged ions instead of electrons solves some old mysteries, and defied old fallacies.
By understanding that the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions) defines the value of the resting potential we defied the decades-old fallacy that some mystic concerted operation of ion channels and some linear combination of purely-defined permeability values define the resting potential of neurons. We calculated by using the principles of classic electricity that neurons handle charge as electricity describes (electrolyte) condensers. We introduced a resting and an excited state and established that that the neuron works as a simple control circuit where the set-point is defined by the construction and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat absorption" by neurons, the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH guessed in the physical picture behind their empirical model and mathematical equations, and introduced the correct physical model by providing also its detailed mathematical description.
We could theoretically underpin (see Fig. 6) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH, although not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation on the membrane, as well as polarization in the electrolyte surrounding the membrane. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Considering the presence of electrolytes on the two sides of the membrane, a correct value for the resting potential can be derived numerically from the first principles. Those layers have a decisive impact on the neuron’s electrical operation. The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [7] of science must be considered.
We have shown that the neuron is a well-designed, simple control circuit with separate ground and excited state mechanisms. The geometry and the overall concentration of the cellular fluids determine the nominal values of the electric and thermal gradients. The ion currents through the permanently open "resting" channels and the gated "transient" ion channels, respectively, deliver ions, which means they affect the gradients until the equilibrium state is kept or restored.