### v2

# The unified cross-disciplinary model of the operation of neurons

Abstract
Physics perfectly describes neuronal operation, provided that we take into account that biology uses slow positive ions instead of electrons as charge carriers and remove untested ad-hoc hypotheses that contradict science’s first principles. We also incorporate recent experimental discoveries into the outdated classic theoretical description. Lipid mechanisms are really very important for cellular biology, but they are certainly not suitable for describing the phenomena we discuss. We introduce the correct physical model, significantly enhancing the classic Hodgkin and Huxley (HH) model; furthermore, the fundamentally bio-electrically triggered operation causes changes in electrical, mechanical, and thermodynamic properties of living matter. We derive the resting potential from first principles of science, showing that it has no relation to an ad-hoc linear combination of mobilities or reversal potentials as the Goldman-Hodgkin-Katz (GHK) equation claimed. Furthermore, we derive an "equivalent thermodynamic electric field" that enables discussion of, among others, the operation of ion channels, their ion selectivity, and voltage sensing. We demonstrate that a simple electrical-thermodynamic control circuit regulates neuronal operation, setting and maintaining a stable resting potential and handling an unstable transient process known as the Action Potential (AP). Its setpoint entirely defines the resting potential, explaining its robustness during growth and evolution. Our cross-disciplinary approach naturally fuses the electrical and mechanical/thermodynamic description of neuronal operation, resolves the decades-old mystery of "heat absorption" and "leakage current" (with their far-reaching consequences), and derives the thermodynamic description of neural computing. We defy that science cannot describe life.
keywords:
slow currents, neuron’s physical model, resting potential, leakage current, negative feedback, thermodynamic driving force, Goldman-Hodgkin-Katz equation, reversible heat production, thermodynamics of neurons
The table of contents is present only for the comfort of reviewers
1 Introduction
The excellent textbook [1] separates the neuron membrane’s states into resting and transient states. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; here we tell the second half. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential"  [1], page 126. In the resting state, two effects are considered. On the one side, the potential is created by a static charge-up: "the cell membrane potential results from the separation of net positive and net negative charges on either side of the membrane". However, as we discuss, it is only a small part of the potential: the polarization of the electrolyte’s dipoles contributes the rest. (For the sake of clarity, both mechanisms are known in physics. The word ’polarization’ is reserved for rearranging charges in multiply-charged systems without the application of an external current. The changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even if physiology refers to it as polarization. External currents contribute to the changes of potential.) On the other hand, it was claimed that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [1], although it was question-marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [2]. The claim is rooted in the hypothesis and the measurements by Hodgkin and Huxley [3], as we discuss in section 2.6.4. Their numeric figure was correct, but they mistakenly interpreted the voltage they measured as being produced by a "leakage current" on the membrane’s parallel resistance. Furthermore, their interpretation leads to causality issues [4]. Paper [5] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways, but did not account for the electric potential. We derive how that electric potential is produced by separating charges and how its value can be calculated from first principles.
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [6]. More precisely, inseparable thermodynamic and electric processes keep the potential set by the inseparable electric and thermodynamic backbone defined by the lipid structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [7].
The book [1] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, as it discusses only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization. We demonstrate in section 3 that classical electric methods enable us to calculate the membrane’s potential resulting from charge separation and polarization. However, one can make some physically plausible assumptions to provide numerical figures. The electric processes are inseparable from the corresponding thermodynamical processes [8], which we discuss by deriving an "equivalent thermodynamical electric field" in section 2.3. The other minor effect is a low-intensity "passive flux" (material transport) produced by the "resting channels". They are ion channels, not necessarily only across the membrane; they also include the ion channels in the Axon Initial Segment (AIS) [9, 10, 11, 12, 13]. The "resting current" is about two orders of magnitude lower [14, 15] than it was assumed at the time of writing the book; correspondingly, its role is much smaller in the transient state, and it even drastically changes the electric model used in producing an AP which we discuss elsewhere [16].
2 Membrane’s electricity
The importance of the subject was correctly estimated decades ago [17]: "For the past forty years, our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of HH. Their work has had a far-reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. …The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system." To fix the flawed and incomplete hypotheses has a similar importance.
It was already stated that the membrane’s resting potential is, of course, of electrostatic origin [18, 19, 20, 2, 21]. The closed volume of the neurons drastically differs from the infinite space: the charges repel each other, but since they experience the counterforce of the limiting surface (the membrane), a static mechanical pressure and a shock wave are also created, that naturally connects the electric and mechanical effects, as well as the electrical, mechanical, and thermodynamic theories of operation. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [8]. Furthermore, the features of living matter may change while being measured in physics laboratories as usual: "the construction is different from anything we have yet tested in the physical laboratory" [22].
In the absence of external influences (which includes the lack of a separating membrane), both gradients are balanced and are at zero. When this state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [8] , and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions (the disciplinary way) leads to incorrect conclusions and theoretical interpretations.
2.1 Condenser
Biology represents a complex case where the phenomena cannot be reasonably reduced to a single discipline, as classic inanimate science does when deriving its laws for "material points", as A. Einstein coined, abstracted to either mass or charge, but not both. Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [23, 24], but does not explain that the negative and positive ends of lipids in an electrolytic solution attract and absorb ions, forming a charged layer on both surfaces of the membrane. In this way, the membrane in electrolytes acts as a condenser. However, this condenser comprises two charged sheets on the surfaces of the membrane with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (as opposed to conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolytes, so its thickness is far from the idealized zero-thickness known from classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electric point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
Cell science does not establish its electrical phenomena; instead, it attempts to elucidate complex lipid activities that explain charge transfer, resting and transient potential, and the entire biological operation using mathematical formulas without a realistic physical background. Electrophysiology combines currents of electrons and ions, and uses Ohm’s Law for its admittedly non-Ohmic systems. It introduces foreign (clamping) currents into biological systems, attributing their effect to a changed conductance of the membrane without explaining how the voltage and current can be independent of charge carriers.
2.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electric forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electric field (which can be of electrical or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electric field fluctuates, so their state is dynamic: the molecules dissociate, and the free ions recombine. Furthermore, the process plays an important role also in the energy supply of the neurons: the Adenozine Triphosphate (ATP) hydrolyzes molecules and the electric field near the membrane moves the ions to the "condenser plates". Depending on their environment, the substances can exist in base, ionized, and polarized states (and behaviors).
In biological cells, in a resting state, a small (approximately $10^{-3}$) portion of the molecules dissociates (i.e., the ions separate from their counterparts with opposite charges), and they move freely within the volume. In other words, the electrolyte liquid can then conduct electricity due to the mobility of the positive and negative ions (of course, if some electric/thermodynamic field provides the required driving force). The rest of the molecules can be more or less polarized, allowing for the possibility of producing an internal electric field in the solution. In a transition state, all components (including ions, molecules, and ATP) have gradients and they move towards their balanced state.
2.3 Deriving the "thermal electric field"
In a segmented electrolyte, the electric charges are typically globally balanced; however, they may become unbalanced locally due to physical reasons. The Nernst-Planck electrodiffusion equation in a balanced state
$\frac{d}{dz}V_{m}(z)=-\frac{RT}{q*F}\frac{1}{C_{k}(z)}\frac{d}{dz}C_{k}(z)$ (1)
describes the mutual dependence of the spatial gradients of the electric and thermodynamic fields on each other. In good textbooks (see, for example, [25], Eq (11.28)), its derivation is exhaustively detailed. In the equation, $z$ is the spatial variable across the direction of the changed invasion parameter, $R$ is the gas constant, $F$ is the Faraday’s constant, $T$ is the temperature, $q$ the valence of the ion, $V_{m}(z)$} the potential, and $C_{k}(z)$} the concentration of the chemical ion. The two sides are the derivatives of the equation
$V_{m}=\frac{RT}{q*F}\ln{\biggl(\frac{C_{k}^{ext}}{C_{k}^{int}}\biggr)}$ (2)
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine additively quantities from the intracellular and extracellular sides, whether concentrations, or mobilities, or Nernst voltages, as it happens in the GHK equation (see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [8]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [26], the two differentials must change in the same direction given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (2) results in a negative potential difference).
It is a common fallacy in physiology that the applied potential $U$ generates a current. Instead, its gradient ($E=U/d$) has the effect, where $d$ is the distance, for example, between the clamping point and the membrane or the AIS. We assume that, in a balanced state, the concentrations on the two sides of the membrane are $C_{k}^{ext}$ and $C_{k}^{in}$, respectively; furthermore we assume that
$\frac{dC_{k}}{dz}\approx\frac{C_{k}^{ext}-C_{k}^{in}}{d}$
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, furthermore, that U=EdU=E*d, we can derive the "equivalent thermodynamical electric field" for the case of a permeable ion channel in the membrane
$E_{thermal}^{C_{k}}=-\frac{RT}{q*F}\frac{1}{C_{k}^{ext}}\frac{C_{k}^{ext}-C_{k}^{in}}{d}\approx-\frac{RT}{q*F}\frac{1}{d}$ (3)
It does not depend on the concentrations, but depends linearly on the temperature and reciprocally on the membrane thickness (given in $nm$); furthermore, it is valid separately for all $C_{k}$
$E_{thermal}^{C_{k}}(d)=-2.585*10^{-2}\frac{1}{d}\quad\biggl[\frac{V}{m}\biggr]$ (4)
Figure 1 depicts the field’s dependence on the membrane’s width and how it compares to the charge-generated electric field. The figure illustrates the setpoint, the boundaries of the changes, and that the interesting phenomena happen within the cross-section of the shaded areas. The figure is an illustration, but the figures are not far from the true ones. For a 5nm5\ nm membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Eq. (18). (Notice that the membrane’s thickness may change during the action potential, see [27]: the rush-in extra charge compresses the membrane. Similarly, the concentration of the electrolyte in proximity to the membrane also changes.) Correspondingly,
$U_{thermal}^{C_{k}}(5)=-25.85\quad\biggl[{mV}\biggr]$ (5)
Figure 1: The thermodynamic "electric field" depends on the membrane’s thickness, see Eq.(4). The electric gradient depends on the charged layer’s concentration and thickness; see Eq.(18). The dots mark measured data shown in Table 1.
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electric field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, joint charge-generated electric fields exist, and they equal, per ion, the concentration-generated "electric fields", see section 3.1.1.
In the steady state, some other forces also contribute to the mentioned forces. To discuss how nature restores the steady state when a microscopic change happens in a balanced state of a biological solution, we write the well-known Nernst-Planck equation (see Eq.(1)) in a slightly extended form:
$0=\underbrace{e_{el}*E_{Gap}^{Total}(c,\Delta z)}_{Electrical~force}+\underbrace{e_{el}*E_{thermal}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{k}}}(d)}_{Thermodynamical~force}\quad\underbrace{\bigl(+F_{ext}(z)\bigr)}_{Constraint}\quad\underbrace{\bigl(+F_{Transp}(z)\bigr)}_{Transport~force}$ (6)
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added an external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [28], with a lot of cited references.)
Expressed differently
[…]
That means when describing an ionic transfer process, we must not separate the electric current from the mass transfer (in other words, to discuss electrical or thermodynamic models): they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electric term is not. To be entirely balanced, the system must be balanced to all elements. This way, changing one concentration implicitly changes all other concentrations and the electric field.
Notice the special role of the constraint force by the closed volume represented by the membrane’s wall. The thermodynamic contribution also means a pressure from the colliding particles, and the electric repulsion exerts a force on the membrane’s wall. Those disciplinary forces act simultaneously and inseparably.
2.4 Fallacies about biological electricity
When discussing balanced states, no transport occurs, so the transport force cancels, and the mobility has no role (when speaking about mobility provided by protein mechanisms, one should derive the corresponding contribution to Eq. (7)). The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions on the opposite side, and the membrane mechanically provides the needed counterforce. This way, no resultant force acts on the ions; the two gradients persist. When the ions can move freely between the segments, they will continue to move until they produce a concentration gradient (a thermodynamic force) for the given ion that counterbalances the electrical gradient (the electrical force), and then the transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq. (27)) in a viscous fluid.
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electric current in solids.
It was a colossal mistake to forget that biological currents comprise ions with their electrostatic forces. It led to fallacies, among others, including leakage current (see Section 2.6.4) and heat absorption (see Section 4.6). Their finite (Stokes-Einstein) speed, see section 2.9, due to the ions’ mass and the different conduction mechanism, explains the spatiotemporal behavior of living matter [29] (contrasted to Newton’s ’instant interaction’).
It was another colossal mistake to introduce equivalent circuits with constant-value voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason, and prevents understanding how the competition between thermodynamic and electric processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 2.9; this way attributing the change of an electric entity to the biological material. This assumption neglects that a driving force is needed to move ions (denies Newton’s laws of motion). Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
The classic textbook [30], appendix A, correctly warned (especially when working with concepts of an electric theory of neuronal operation): "An intuitive grasp of concepts of electricity and electrical circuits is helpful for understanding some of the basic theory in cellular neurophysiology. …Under ideal circumstances, the physical act of measuring a neurophysiological event would have no effect on the electrical signal of interest. Unfortunately, this is seldom the case in neurophysiology."
One of the bad examples is grasping the electrical characteristics by clamping (we are not speaking against the method, but emphasizing the need for its correct interpretation). The textbook also provided the correct interpretation of the measured quantities: "In describing the resistance (or Ohm’s law) properties of neurons, one either injects current and measures the resulting change in voltage (a current clamp) or changes the voltage to different values and measures the current necessary to hold (or clamp) the voltage to these values (a voltage clamp)."
As we discuss in section 4.1, physiology does not realize that in resting and transient states, different physical processes take place and correspondingly, different circuit models shall be used. In the resting state, the low intensity ion current is "shunted" by the ion channels in the membrane’s wall; moreover, the "parallel RCRC" model works sufficiently well. Since the current is slow, it is not simply distributed, but only the "overflow" reaches the AIS. However, as the current intensity exceeds the ion transmission capacity of the "resting ion channels", the "transient ion channels" play an increasing role. Above the voltage threshold, the serial $RC$ circuit is the correct model. Until the neuron’s voltage reaches its threshold, a "parallel" and a "serial" resistor work simultaneously, in this way implementing a mixed model. Measuring the "input resistance" of neurons actually measures, in this sense, a mixture of the "input" and "output" resistance corresponding to the two models, given that the current is distributed between the one through the membrane and the other through the AIS; plus the measuring electrode represents one more drain.
At the beginning, it was clearly told [31], but forgotten: "Since it is quite generally believed that the depolarization of a nerve fiber membrane, during excitation and propagation, involves an increased permeability to ions there have been many attempts to detect and measure this change as an increase in the electrical conductivity. … In these cases, the measuring current was also the stimulating current, and it was not possible to analyze the changes satisfactorily." Moreover, during the passage of an impulse, a "charge wave" passes along the axon and the changed number of charge carriers in that cross-section, see Eq.(20), changes the current that creates the illusion that the conductance at the measured cross section changed. In our terminology, the stabilizing feedback voltage provides an external force, and it also contributes an external current. The feedback amplifier’s electron current (converted to ion current) is provided externally and is added to the closed system, without including the measurement device into the closed system. That interpretation of a closed system denies the fundamental symmetries of science (among others, conservation of charge, mass, impulse, and energy). Moreover, it significantly contributes to the claim that "laws of science cannot describe life".
All conservation laws are correct and valid, only "the construction is different" [22] and the "test method" is wrong. Not mentioning that under some conditions, the "system under test" also "produces" charge carriers by opening ion channels. The "construction" needs revisiting concepts of electricity for living matter instead of projecting concepts valid for inanimate matter only, to living matter. That wrong "grasp of concepts of electricity" suggests that Ohm’s Law is not valid for living systems. The correct statement is that Ohm’s Law is not valid for arbitrarily chosen currents and voltages. Without that revisiting, by saying that living systems show non-Ohmic behavior, biology claims that it applies good laws inappropriately to a good system. The consequences are far-reaching: physiology uses magic changing conductances in circuits with fixed-voltage batteries, that blocks not only understanding the thermodynamic and mechanical effects (and introducing the corresponding theories [32, 33, 34, 35, 36, 27, 37, 38, 6, 39, 40, 41]), but even consistently describing the electric phenomena. We are not speaking against using the concept of conductance, but rather against deriving its value incorrectly by including foreign current compositions, which obscures the actual electrical behavior of biology. In addition, by introducing current feedback (which, by definition, compensates for any gradient in the system), physiology applies an opposite-phase control unit against the neuron’s native control unit and concludes that, after compensation, there is no gradient in the system; in other words, that life exists without needing a driving force.
Physiology is not aware that, inside the neuron, there is a condenser with a voltage difference between its plates. Instead, it believes that a ‘leakage current’ flows through resistors distributed over the membrane. That current is special in that, among other things, it has neither a chemical composition nor charge carriers; furthermore, it can not only emit but also absorb heat. Neuroscience observes the thermodynamic and mechanical changes (for a comprehensive review, see [33]) but cannot connect them to the electrical phenomena. Lacking the required physical background, biology claims that physical laws cannot describe processes in living matter, without providing a plausible and self-consistent theoretical model.
2.5 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electric effects connected to membranes and nanopores [19] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
In the calculations below, we require the ’surface charge density’ (interpreted as an ’infinitely thin layer’ in physics, expressed in $C*m^{-2}$). We can derive (assuming singly-charged ions) the volume charge density
$\sigma_{V}(c)=c*N_{A}*e_{el}=c*6.023*10^{20}*1.602*10^{-19}=96.4*c\quad\biggl[\frac{C}{mM*m^{3}}\biggr]$ (8)
where $N_{A}$ is the number of ions in a millimolmillimol and $e_{el}$ is the elementary charge. Notice that all ions contribute to the charge density, so simply $c=\sum_{k}C_{k}$.
Here, we run into a conflict between the ’infinitely thin’ layer of physics and the biologically implemented (finitely) ’thin physical layer’, so we use a thickness parameter $\Delta z$. By assuming an arbitrary ’physical layer thickness’ $\Delta z$, we can calculate the surface charge density
$\sigma_{A}(c,\Delta z)=96.5*c*\Delta z\quad\biggl[\frac{C}{m^{2}}\frac{1}{mM*nm}\biggr]$ (9)
We interpret classical physics’ notions from the theory of ’continuous’ electricity, interpreted for ’infinite’ cases, for the finite world of biological objects and the quantized atomic world. Our idea is similar to the one used by Boltzmann for connecting the continuous medium to quantized particles. When ions are present in an electrolyte having concentration $c$ in a ’conducting layer’ of thickness $\Delta z$ on an isolating surface, they produce a field
$E_{z}(c,\Delta z)=\frac{\sigma_{A}(c,\Delta z)}{\epsilon_{o}}=10.89*10^{3}*c*\Delta z\quad\biggl[\frac{V}{m}\frac{1}{nm*mM}\biggr]$ (10)
Figure 2: The electric field of a disc of charge can be found by superposing the point charge fields of infinitesimal charge elements. This field can be facilitated by summing the fields of charged rings. “From Hyperphysics by Rod Nave, Georgia State University". ©hyperphysics.phy-astr.gsu.edu/hbase/
2.6 Two segments
Now, we separate the volume into two segments and compose the segments from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electric field’s $z$ component of those plates in the points as shown in Fig. 2. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electric field is zero inside the conducting layer), so the electric field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into the ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [42]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electric fields. Although not explicitly stated, we consider electric double layers to be present near the membrane, and ions from the dispersion medium are adsorbed on the particle’s surface, depending on the chemical properties of the medium.
Correspondingly, the theory non-equivocally explains the potential near the membrane surface; see the discussion and review in [43], mainly their Figure 1. That model considers a "symmetric" condenser, that is, both plates contain charges of the same sign. The Donnan function is a (discontinuous) step-like potential. Its continuous version introduces a mathematically smooth, non-physical function at the boundary of the charged layer. One must describe a physically discontinuous layer (electrolyte, ion layer on the membrane’s surface, and isolating membrane) with a single continuous function and gradients (potential and concentration) with physical meaning.
The behavior near the boundary layers is problematic, primarily because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electric field; the charged layer and the electrolyte jointly produce the electric field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electric field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their electric force is a fragment of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it generates an electric field similar to that of a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electric field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separates the segments). The membrane does not need to generate any counterforce (except for collisions).
2.6.1 Charge layers in segments
As we show, the finite thickness will result in a lack of balance (introduces inhomogeneity and creates a voltage and concentration gradient) proximal to the surfaces of the membrane, even if the concentrations on the two sides are the same. Changing the bulk concentration or potential in one of the segments creates a corresponding gradient across the separating membrane, increasing the inhomogeneity proximal to the membrane. The ions will experience an extra force due to the gradient; however, the mechanical counterforce of the membrane will keep them back in the segment. The concentration and potential, inseparably and having the same time course, will change across the two sides of the membrane, just because of the gap’s physical features the membrane represents.
We consider that the segment is composed of electrically conducting discs (the ions are free to move on the surface), and the charged discs’ contribution to the electric field at point $P$ (see Fig.  2) can be calculated as known from the theory of electricity. Due to the symmetry in the $z$ direction, in a homogeneous solution, the resulting electric field in the plane perpendicular to the $z$ direction is zero, as we have contributions of equal magnitude with opposite signs.
Notice that here we used the abstraction of an infinitely thin "charged sheet", and there is a step-like gradient in the electric field in the gap. The physical reality is that the charge is represented by ions, which simultaneously represent a mass. In equilibrium, the forces due to the voltage and concentration gradients must be counterbalanced by an external force. An interference between science disciplines can also manifest here. We know simultaneously that at the boundaries of electrolytes, different interfaces, including electrically neutral electric double layers, can be formed by only partly known processes [44]. The presence of those structures makes drawing quantitative conclusions hard.
2.6.2 Simple conducting sheet (a physical condenser)
First, we consider a condenser that obeys the laws of classical electricity and has the geometrical size of a biological neuron. We apply the physics terms to that pseudo-biological object and test whether the values we derive are consistent with physiological phenomena. A tiny (according to [30], page 12, about $2*10^{-3}$) portion of the positive ions leaves their negative counterions behind and forms a thin positively charged ion layer on the surface of the membrane (of course, the same happens on the opposite side with negative ions). Experience shows that – also in biology – two very thin layers of ions are formed on the two sides of the membrane of finite width, see the caption of figure (Fig. 11.22 in [45]): "The ions that give rise to the membrane potential lie in a thin (<1nm<1\ nm) surface layer close to the membrane" (As shown below, these ions give rise only to part of the potential, and the dipoles in the electrolyte contribute the rest). A sub-nanometer-thick conducting layer covers the two surfaces on the top of a 5nm5~nm thick insulator. So, we model the cell with two finite-width "conductive sheet" layers. These layers represent an electric condenser (so we can calculate the internal electric field between the plates) and counterbalance each other’s electric field outside the condenser. We may assume that the solid surface represents the needed counterforce to keep the charges at rest.
We cannot expect geometrically flat surfaces, as some biological objects of up to 1nm1~nm in size are present on the surface. So, we must assume a physical ’uniformly charged sheet’ with an average thickness $\Delta z=0.5~nm$ on the surface. For the sake of simplicity, we assume a step-like function for the electric field. It is zero in the segment outside the "conducting sheets" (the ’bulk’ portions) and inside the sheets. Between the sheets, it jumps to the value of the electric field of a charged sheet with surface density $\sigma_{A}$ (see Eq.(9)). (In this picture, we see a jump in the value of the electric field on the two sides of the physical layer. We assume that the physical layer somehow emulates the well-defined boundary on the side opposite to that proximal to the membrane. Here, we still assume that a mechanical counterforce due to the surface’s roughness keeps the charges inside the charged layer.)
We assume that all unbalanced (dissociated) ions are in the layer, and the rest contains no dissociated ions; furthermore, the concentration of the ions is unchanged in the segment. In the case of a 400mM400~mM solution and a 0.5nm0.5~nm ’thin physical layer’ Eq.(10) evaluates to an electric field
$E^{Gap}_{Classic}(400,0.5)=0.22*10^{7}\biggl[\frac{V}{m}\biggr]$
and at a 5nm5\ nm membrane thickness, the voltage across the plates becomes
$U^{Gap}_{Classic}(400,0.5,5)=E^{Gap}_{Classic}(400,0.5)*10^{-3}*d=10.84\ \bigl[mV\bigr]$ (11)
Our simple model seems to be a strong oversimplification. The number of uncompensated ions needed for the cell, using the method in [30], is $0.88*10^{7}$. The calculated values are about a factor of $5$ lower than the experimentally derived values. "An electrical potential difference about 50100mV50-100\ mV … exists across a plasma membrane only about 5nm5\ nm thick, so that the resulting voltage gradient is about 100, [^0] V/cm100,000\ V/cm[45]. "The number of uncompensated ions needed for the cell is $4.7*10^{7}$[30]. We are in the correct order of magnitude, but we arbitrarily assumed a layer thickness, a non-dielectric medium, and a physically unreasonable step-like electric field. The deviation from the experienced values suggests that dipoles in the electrolyte bulk significantly change the achievable electric field and potential across the plates.
Figure 3: The oppositely charged parallel surfaces of the membrane are treated like conducting plates (infinite planes, neglecting fringing), then one can use Gauss’s law to calculate the electric field between the plates. Presuming the plates are at equilibrium with zero electric fields inside the conductors, the result from a charged conducting surface can be used. “from Hyperphysics by Rod Nave, Georgia State University". ©http://hyperphysics.phy-astr.gsu.edu/hbase/
A usual parallel plate condenser (shown in Fig. 3) can be derived with such a model. We have two infinite conducting sheets and an insulator layer between them. The magnitude of the opposite charges on the two plates must be the same as in the classical picture. Given that the amount of surface charge densities is concluded from the ionic concentration in the segment (in a balanced state), the sum of ionic concentrations must be the same in the segments. A constant electric field is present inside the membrane (across the plates of the condenser), and there is zero electric field inside the parallel conducting plates and outside the condenser. In this ideal picture, the charges are aligned on the border of two (infinitely thin) conducting layers and cannot move. The attractive force between them and the opposite charges on the other plate keeps them fixed in the direction of $z$, and the repulsive force between the charges with the same sign, the surface density in the $(x,y)$ plane remains uniform: the infinitely thin plates are equipotential. This picture illustrates the equilibrium state of charges, consisting of infinitely small non-dissociating charge carriers and perfectly smooth surfaces.
2.6.3 Condenser in dielectric material
In its integral form, Gauss’s flux theorem states that the flux of the electric field out of an arbitrary closed surface is proportional to the electric charge enclosed by the surface, irrespective of how that charge is distributed. The surface layer represents a steep potential and concentration gradient. Above, we assumed that the counterforce that keeps the charges in their place against their electric attraction on the membrane’s surface is a mechanical force: the ions cannot pass through the membrane. However, such a counterforce does not exist on the side toward the ’bulk’ part of the segment. The charges on the plates do not generate an electric field toward the bulk, but the concentration they represent does, according to the Nernst-Planck equation, provided there are charge carriers in the bulk. We hypothesize that virtual electric charges exist in the electrolytes, and their field provides the missing electric field.
When explaining the effect of dielectricity, we must explicitly consider the duality of ions that they obey laws of electricity and thermodynamics simultaneously; furthermore, the complexity of the electric structure of the solution and that the charge carriers have finite size (see "electron size" vs. "dipole size"). According to the theory of electricity, the free charge carriers (the dissociated ions) are located on the surfaces where they form powerfully charged thin layers (condenser plates) in the segments separated by a membrane on the two surfaces of the membrane; two proximal charge layers on the surfaces of the membrane. We assume that those ions behave as point charges, i.e., due to the attraction from the opposite charges on the opposite plate, they do not produce an electric field toward the side of the bulk of the electrolyte layers. In the classical picture, those layers represent step-like gradients in the electric field along the $z$ axis, and the constraint that the ions must not enter the membrane provides the needed counterforce.
From a thermodynamic point of view, a driving force acts as long as a concentration gradient between neighboring layers exists. From an electrical point of view, the layers are simple parallel-plate condensers that produce no electric field outside their closed volume and are connected serially; plus, the top layer has free ions. The electric field is proportional to the bulk concentration in the layer. Changes in the local electrical gradient may also alter the degree of dissociation and polarization, producing a graded local electric field. These two driving forces have opposite directions and in a balanced state, the same magnitude. The counterforce acts in a way that constrains the molecules to stay in their layers; it adapts (also compensates for the different interaction speeds) while the gradients are changing. It presses the proximal layer to the membrane and the neighboring layers to each other.
Due to the membrane’s finite width and surface roughness, the conducting layer has a finite width. The electric field is homogeneous within the layers. The rest of that electric field toward the bulk layer directs the dipoles proximal to the neighboring dipole layer, and their polarization increases; in other words, they "produce" an electric field. The final effect is that a dipole layer is attracted to the charged layer, and it forms another layer that shows a charged layer toward the bulk. The process repeats and results in a decaying electric field based on the distance from the membrane’s surface.
In our model, we build the volume of the electrolyte from thin layers of directed dipole molecules (the size of molecules limits the thinness $\Delta z$) in the volume, separated into two segments with electrolytes by a membrane with a finite thickness $d$ (that is, a point’s distance in the electrolyte from the two disks will be $z$ and $z+d$, respectively). The contributing potential of an infinitesimal volume on the axis at point $z$ due to the charged sheet is
$dU(c,\Delta z,z)=\underbrace{E^{Gap}_{Classic}(c,\Delta z)*z}_{Potential~from~sheet}*\frac{1}{z^{2}}=E^{Gap}_{Classic}(c,\Delta z)\frac{1}{z}$
The total potential due to all elements $dz$ from the left side is (in the mathematical formulas below, we express the distance in units of $\frac{z}{d}$, so here $z$ is dimensionless).
$U_{left}(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d\int_{-\infty}^{0}\frac{1}{z}dz=\\ E^{Gap}_{Classic}(c,\Delta z)*d\biggl[\ln|z|\biggr]_{0}^{\infty}$ (12)
In the case of a single-segment volume, within the segment, a similar potential with an opposite sign is generated by the charges on the neighboring sides of the considered charged sheet, counterbalances the potential described by Eq.(12). However, when a membrane with thickness $d$ separates the segments (with no charges in the gap), the potential on the right side will be
$U_{right}(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl[\ln|z+1|\biggr]_{1}^{\infty}$ (13)
We can write the electric field in the form
$E(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*ln\biggl(\frac{z}{z+1}\biggr)$ (14)
That is, the gap sets the potential difference across the membrane to
$U(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl(\biggl[\ln|z|\biggr]_{-\infty}^{0}-\biggl[\ln|z+1|\biggr]_{1}^{\infty}\biggr)$ (15)
We use the approximation that $\ln(\infty)\approx\ln(1+\infty)$, and we arrive at that
$U(c,\Delta z,d)=E^{Gap}_{Classic}(c,\Delta z)*d*\biggl[ln\biggl(\frac{z}{z+1}\biggr)\biggr]_{0}^{1}$ (16)
describes the potential across the plates; that is, the classic potential shall be multiplied by the result of the integral. It was experimentally confirmed [46] that the potential of the membrane is proportional to the logarithm of the distance from the membrane.
2.6.4 Leakage current
Although through $\Delta z$, the ion’s water-related behavior can influence the surface charge density, one can conclude that when only one type of ion is solved in the liquid, approximately the same concentration difference between the segments produces approximately the same "thermodynamic electric field" that can counterbalance the electric field of the biological condenser.
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage could be generated; see Fig. 3 in [15]: the leakage current is insignificant. More precisely: low intensity slow current flows between nearby ion channels inside the membrane, without involving the AIS. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a 42.5mV42.5\ mV voltage [3]. In HH’s picture, the question what ion constitutes the current, remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without a driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leakage^{+/-}$ ions (??) (depending on the direction, it could be positive or negative), one generates voltage while the other not, and neither of them changes the concentration neither on the departure nor on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law. BTW: [47] discusses the operation of the membrane, in some modified HH approach: it replaces "leakage" current with "pump" currents. In the wrong model, such a modification does not matter.
The idea of leakage current was wrong: the voltage HH measured [3] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on the two sides of the membrane, even when no current flows. Measuring the energy consumption of neuronal operation [14] confirms that no resting current, in the sense as HH used the concept, exists. The existence of leakage current would be against, among others, the overall metabolic efficiency of neuronal operation.
2.7 Bridging the micro- and macro views
We assumed that the function can be interpreted for regions $(-\infty,0)$ and $(1,+\infty)$. However, we have arrived at the boundary between quantized and continuous electricity. We use a trick similar to the one Boltzmann used in his famous equation, except that we calculate the number of charge carriers from geometrical instead of statistical assumptions. The minimal thickness of layers is limited, at least by the size of the ion. Furthermore, the membrane’s surface is not flat; there are structures (mainly lipids) with sizes of up to 1nm1~nm, so it is likely realistic to consider a layer thickness of 0.5nm0.5~nm as we did above (on our mathematical scale $0.1=\frac{0.5~nm}{5~nm}$), for which a multiplier of $1.7$ between the classic and the electrolytic fields exists. A similar calculation using a layer thickness of 0.1nm0.1~nm (on our mathematical scale, 0.02) results in a multiplier of $3.2$; so we use a multiplier of value 3 (that corresponds to a layer thickness of 0.125nm0.125~nm (the radius of a $Na^{+}$ ion is 0.116nm0.116~nm)). Correspondingly, we assume an electric field
$E^{Gap}_{Dielectric}(c,\Delta z)=3*E^{Gap}_{Classic}(c,\Delta z)\quad\biggl[\frac{V}{m}\frac{1}{nm*mM}\biggr]$ (17)
due to the dielectric properties of the segment. This contribution is due to the dielectricity in the segment (for the sake of simplicity, we have neglected the relative permittivity). It is to be added to the value of the classic contribution, the field generated by the conducting plates, so the final multiplier is $4$. (The experience is known in technical electricity: the electrolytic capacitors achieve several times higher charge storage capacity by using (pseudocapacitance). By using "roughened anode foil", the thickness of the electrolyte is increased, and the roughening provides the necessary mechanical support. A thicker electrolyte layer wraps the condenser plates, and the dipoles in the thicker electrolyte provide an additional charge storage facility.)
Correspondingly,
$E_{Gap}^{Total}(400,0.5)=4*E^{Gap}_{Classic}(400,0.5)=8.8*10^{6}\bigl[\frac{V}{m}\bigr]$ (18)
and at a 5nm5\ nm membrane thickness, the voltage across the plates becomes
$U_{Gap}^{Total}(400,0.5,5)=E_{Gap}^{Total}(400,0.5)*(5*{10^{-9}})=45\ \bigl[mV\bigr]$ (19)
"The plasma membrane of all cells, including nerve cells, is approximately 6 to 8 nm thick and consists of a mosaic of lipids and proteins." [1], page 71. Using such a value for the membrane’s thickness may result in a value that is up to 60% higher. Furthermore, the usual concentration is also up to 10% higher. Hodgkin, in 1964, measured molarity values in squid axons for ions $K^{+}$, $Na^{+}$, and $Cl^{-}$, (400,50,40-150) inside and (20,440,560) outside, and they provided potential values 5575mV55-75~mV [30]. Given that we used a plausible but ad hoc "charged layer thickness" and gap distance, we cannot expect a better agreement. However, our qualitative conclusions remain valid and can be adjusted quantitatively to the results of dedicated measurements.
Figure 4: The neuronal membrane’s electric field in the function of the distance from the membrane’s surface. The thickness of the ’atomic layers’ proximal to the membrane’s surfaces are also shown. The electric field is linear inside the uniformly charged atomic ion layer on the surface and is described by Eq.(14) in the electrolyte. For visibility, a layer thickness 0.5nm0.5\ nm is assumed. For illustration, a simple capped ion channel in the membrane’s wall is also displayed.
We have the boundary conditions that we know the electric field at the membrane’s surface through the volume charge density, and the Nernst-Planck equation delivers the change of the electric field due to the change in concentration. The physical constraint is that the layer thickness cannot be infinitely small. If we use a parameter for layer thickness equal to that of the surface layer, the calculated electric field distribution will be scalable.
2.8 The dynamic layers
This layer’s importance and operation are in line with the Earth’s atmosphere. Its features drastically deviate from the features of the bulks on its two sides. It is separated by a sharp contour on one side and an ill-defined border on the other; its volume is far from being homogeneous. Gravity keeps it in place, and it is at rest. However, sometimes, for some periods, also other (thermodynamical and electrical) forces are evoked inside it and lead to transient changes, moving huge masses with high speeds inside it. Their thickness and mass are negligible compared to those of the bulks on their two sides, and we can describe the bulks without considering their density, mass, size, etc. Still, this thin layer is responsible for the weather. It can temporarily absorb products of slow processes (water evaporation) and deliver masses of high density (much above its density, such as water, sand, etc.) to continental distances, creating the illusion that it stores that matter. Minor changes (natural ones, such as a slight difference in air temperature, and artificial ones, such as injecting condensation nuclei in clouds) can result in enormous changes. We can even imagine volcanic eruptions as semipermeable gates for material with apparently random operation and distribution of the injected material.
To describe those complex and continuous phenomena at least approximately, we must separate them into stages. We can describe the stages approximately using omissions, approximations, and abstractions, usually considering only one dominant phenomenon. The described phenomena are interrelated in a very complex way and depend on different parameters. To some point, we can describe that thin layer using a static picture and provide an empirical description of its processes, even though we can give some limited validity mathematical descriptions for those stages. However, we understand that for describing the time course of the transition (contrasting with step-like stage changes) between those well-defined stages of the atmosphere, we need a dynamic description to discover the laws of motion governing the processes.
Similar is the case with the neuronal membranes and the neuronal operation. Now, we are at the point where their decades-old static description is insufficient. We need to derive the corresponding laws of motion to describe the neuron’s dynamic behavior. We need a meticulous and unusual analysis to derive them. In a neuron, in the abstraction science uses, we put together only an ionic solution, a semipermeable membrane, and currents that enter and leave it. All these belong to non-living matter. As experienced, at some combination of their parameters and gradients, qualitatively different phenomena happen, which, in the abstraction biology uses, are called signs of life; our system starts to belong to living matter. Given that the approximations, the derived abstractions, and the mathematical formalisms describing them are different for the two cases, it looks like we have two different, only loosely bound worlds. We realize we have arrived at the boundary of non-living and living matters, and we must go back to the first principles of science to clarify where their boundary is. However, by using our approach, we may defy that "the emergence of life cannot be predicted by the laws of physics" [48]. Our artificial duck looks, quacks, and swims like a duck.
2.9 Ionic currents in a cell
On a microscopic scale, a charge generates a potential field that acts on other charges. It is a common fallacy in physiology that a "potential field travels": neither potential nor current changes without a physical movement of charges. In a conducting wire filled with ions (the solids have a different current transfer mechanism), there are free charges; their number per unit volume is given by $n$, and $q$ is the amount of charge on each carrier. If the conductor has a cross-section of $A$, in the length $dx$ of the wire, we have charge dQ=qnAdzdQ=q*n*A*dz. If the charges move with a macroscopic speed $v=\frac{dz}{dt}$, at the macroscopic level, we define the current $I$ as the charge moved per unit of time as
$I=\frac{dQ}{dt}=q*n*A*v$ (20)
The effect was hypothesized at the very beginning: "it seems difficult to escape the conclusion that the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment." "it is necessary to suppose that there are more carriers and that they react or move more slowly" [49]. The charged ions, attracted by the opposite charges on the other plate and repelled toward the single exit point AIS, really move slowly and behave as if they were a component of the membrane.
Notice that if any of the factors is zero, the macroscopic current $I$ is zero. Microscopic carriers must be present in the volume, and have charge, the cross section must not be zero, and the charge carriers must move with a potential-assisted speed, which needs an electric field generated by the system, see Eq.(7). When there is no constraint force, the thermodynamic and electrical forces will result in an effective force moving the ions. That is, the ions can move without an external electric field, as experienced in biological systems. One of the fundamental mistakes by HH was neglecting the Coulomb force for ions’ electric interaction, and that the electrical potential is created by diffusion processes instead of ideal electric batteries.
In biological systems, where ions represent the "current", it is either a native current (without an external potential) or an artificially injected current or potential generating a current. This way, the current is consistently accompanied by a change in concentration gradient, as the moving ions represent both mass transfer and charge transfer simultaneously. The potential and current are connected through the features of the medium (material) that hosts our measurement.
According to Stokes’ Law, to move a spherical object with radius $a$ in a fluid having dynamic viscosity $\eta$, we need a force
$F_{d}=6*\pi*\eta*a*v$ (21)
(drag force) acting on it. A (microscopic) electric force field $\frac{dV}{dz}$ inside the wire would accelerate the charge carriers continuously
$F_{electric}=k*E(z)*q$ (22)
with a constant speed $v$. It is not the drift speed because of the electric repulsion, it is a potential-assisted or potential-accelerated speed that can be by orders of magnitude higher. The medium in which the charge moves shows a (macroscopic, speed-dependent) counterforce $F_{d}$, which in steady state equals $F_{electric}$, that is :
$I=\frac{k*q^{2}*A}{6*\pi*\eta*a}*n*E(z)$ (23)
We extend the original idea by using
$E(z)=E_{electric}+E_{thermal}+E_{external}\quad in\ electrolytes$
The amount of current in a wire is not only influenced by the electric force field (specific resistance) but also by the number of charge carriers $n$. While the latter is commonly considered constant and part of the former, this is not necessarily the case for biological systems with electrically active structures inside. The medium’s internal structure introduces significant modifications (according to E. Schrödinger: "because the construction is different from anything we have yet tested in the physical laboratory" [22]). Applying an additional electric field to a wire can generate varying amounts of current as the number of charge carriers changes due to biological effects. For axons, we use a single-degree-of-freedom system, a viscous damping model, so the ions will move with a field-dependent constant velocity in the electric space; so it takes time while they appear on the membrane. The activity of potential-controlled ion channels in its wall may change $n$ in various ways; furthermore, that change can result in ’delayed’ currents during measurement, for example, in clamping, as physiological measurements witness.
If we have a concentration $C(z)$, in the volume AdzA*dz, we have dQ=C(z)AdxqdQ=C(z)*A*dx*q charge, resulting in another expression for the current
$I=\frac{C_{k}(z)*A*dz*q}{dt}=C_{k}(z)*A*q*v$ (24)
By combining equations (23) and (24):
$v(z)=\frac{k*q}{C_{k}(z)*\eta*6*\pi*a}*E(z)$ (25)
The higher the resulting space derivative (gradient) and the fewer ions that can share the task of providing a current, the higher the speed. We hypothesized (it needs a detailed simulation) that in the case of this charged fluid, the electric repulsion plays the role of ’viscosity’ [8]. The higher the charge density, the stronger the force equalizing the potential; so $\eta$ is the lower, the higher the charge density (proportional to $C_{k}$). For the sake of simplicity, we assume that the speed is proportional to the space gradient of the local concentration and potential. Recall that our equations refer to local gradients only. The electric gradient can propagate only with the speed of the concentration gradient, given that only the chemically moved ions can mediate the electric field [8]. The lower interaction speed limits the other interaction speed if the interactions generate each other.
The dependence of the diffusion coefficient on the viscosity can be modeled by the Stokes-Einstein relation:
$D=\frac{k*T}{6*\pi*\eta*a}$ (26)
so we can express the speed with the diffusion coefficient $D$ and temperature $T$
$v(z)=\frac{D}{T}*\frac{q}{C_{k}(z)}*E(z)$ (27)
It is important to remember for alternating current experiments: the different ions will move with different speeds under the effect of the same $E(z)$, which is the function of the local concentrations. The conduction speed sensitively depends on the temperature, and through it, the shape parameters [50] of the AP and especially the length of the "relative refractory" period [51].
2.10 Imitating slow currents
At the dawn of finding methods for describing neuronal operation, HH published high-precision measurements [49] enabling detailed testing of theories explaining the observed physiological behavior. Their good physical model that “movement of any charged particle in the membrane should contribute to the total current” lacked considering the mutual repulsion and finite speed of those particles (they are about 50,000 times heavier and a million times slower than the electrons and they move in a closed space segment). Furthermore, they have started from the commonly used wrong assumption that conductance is a primary electric entity. They could "find equations which describe the conductances with reasonable accuracy and are sufficiently simple for theoretical calculation of the AP and refractory period". The "reasonable accuracy" and the need for a simple calculation method (performed on a mechanical calculator) obscured that the laws for electrons are not the same as the laws derived for ions.
As discussed in connection with Eq. (20), the current depends linearly on the number of charge carriers $n$. The physical processes changing the number of charge carriers also define the intensity of the currents, providing a way to imitate slow currents by fast currents (by using current generators using special function shapes), where the spatiotemporal time course of the physical process of producing ions in the system modulates the number of charge carriers.
As an example, we discuss the case of slow current traveling in an axon. The ions travel a finite distance with a finite speed, so there must be a delay between their entry and exit times. The case of switching a clamping voltage on, to some measure, is analogous to the arrival of a spike. Initially, the axon contains no ions. In the classical picture, the axonal current flows into the membrane with capacity $C_{m}$ and increases the membrane’s voltage $V_{m}$ with a time constant
$\frac{dV_{m}}{dt}=-\frac{1}{C_{m}}I_{axon};\ V_{m}(t)=\frac{I_{wall}*(1-e^{-\alpha*t})}{C_{m}}$ (28)
that generates a change in the membrane current
$\frac{dI_{m}}{dt}=\frac{1}{R_{m}}\frac{dV_{m}}{dt};\ I_{m}^{on}(t)=g_{m}(V)V_{m}(t)$ (29)
where $g_{m}=\frac{1}{R_{m}}$ is the conductance of the membrane. That is, the measurable current equals the product of the conductance and the clamping voltage. Equs.(3)-(5) in [49] express this relation. If we assume that the axonal current is “fast”, we arrive at the wrong conclusion that the conductance is voltage- or time-dependent. In contrast, if we assume that the axonal current is “slow”, we naturally conclude that Ohm’s Law is correct and valid also for biology: the conductance/resistance is constant, but the number of charge carriers changes, see section 2.9.
There is no voltage-dependent conductance [52]. Instead, the finite speed of ions and the wrong assumption that conductance is a primary entity misleads physiological research. With wording that "conductance changes", one states that charge carriers appear/disappear/reappear; that is, the charge conservation as well as mass conservation are not fulfilled. Claims such as "the current cannot disappear, it has to go somewhere" [25], page 9, are wrong: there is no conservation law for current, which is the derivative of charge with respect to time. The correct physics background behind of the early phenomenon [31] is that the number of charge carriers, at the cross section where the measurement is carried out, changes as wave of ions ions with variable intensity moves slowly along the axon (they are apparently “created” for the static measurement arrangement). The changing conductance is a fallacy stemming from the idea of "fast current" which is valid for electrons but not for slow ions.
In contrast, when the clamping voltage is switched off, the axon is still filled with charge carriers. The resting potential reaches the end at the membrane “instantly” (in this situation, the conduction mechanism is similar to that in metals). The driving force disappears, the ion stream stops, and the ions discharge into the membrane. The lack and the presence of ions in the axon when switching clamping on and off, respectively, produce the difference that “conductances [actually, currents] increase with a delay when the axon is depolarized but fall with no appreciable inflexion when it is repolarized” [49]. The potential is equalized by the AIS current, producing a net exponential decay:
$I_{m}^{off}=I_{Wall}*e^{(-\frac{\alpha}{R_{m}C_{m}}*t)}$ (30)
During the regular operation of a neuronal membrane, after opening the ion channels, a vast amount of ions flows into the intracellular space from the extracellular space, imitating the effect of switching a clamping voltage. The essential difference is that the ions arrive through the axon to the joining point in clamping. In contrast, through the membrane’s ion channels, they directly contribute on the membrane’s entire surface. The membrane’s size is finite, so with a finite current speed, it takes time until the charges on the membrane’s surface arrive at the AIS, as we discussed for the axonal current. Fig. 5 shows the measured rising and falling edges of the clamping experiment performed by HH [49]. The continuous lines show the diagram lines fitted to the experimental data (notice again that the fitted polynomial significantly differs from the correct exponential function around the time 0.5 msec); the dashed line shows the expected result of the rising edge, calculated with the time constant of the falling edge. The number of charge carriers follows a saturating curve, that decreases the current in the axon. HH measured that their time constants differ far outside their experimental error, that would not be possible if they were the result of charge-up and discharge of the same oscillator circuit. The measured rising edge corresponds to a charge-up process where the charge-up current is a saturating curve due to the change in the number of charge carriers in the axon; see Eq.(23).
Figure 5: The measured current, instead of conductance, (Fig. 2 from [49]), demonstrating that two different physical effects result the rising and falling edge of the measured current.
The slow input currents (although due to slightly differing physical reasons for the different current types) are described by the following analytic form valid for fast currents:
$I_{in}=I_{o}*(1-\exp(-\frac{1}{\alpha}*t))*exp(-\frac{1}{\beta}*t)$ (31)
Figure 6: The PSP shape measured by [53] and fitted by function forms given by Eq.(31).
where $I_{o}$ is the current amplitude per input, and $\alpha$ and $\beta$ are timing constants for current in- and outflow, respectively. They are (due to geometrical reasons) approximately similar for the synaptic inputs and differ from the constants valid for the rush-in current. To implement such an analog circuit with a conventional electronic circuit, voltage-gated current injectors (with time constants around 1 msec) are needed. For example, see in Fig 6 the PSP measured function shape published in [53], fitted with our theoretical function Eq.(31). The corresponding voltage derivative is as follows:
$\frac{dV_{in}}{dt}=\frac{I_{o}}{C}*\biggl(\frac{1}{\alpha}*exp(-\frac{1}{\alpha}*t-\frac{1}{\beta}*t)-\frac{1}{\beta}*exp(-\frac{1}{\beta}*t)*exp(1-exp(-\frac{1}{\alpha}*t))\biggr)$ (32)
2.11 Effects of membrane’s electricity
2.11.1 Maxwell demon
Nature builds ’Maxwell-demon’-like objects into the separating membrane: gated ion channels, see Fig. (4). The ion channels operate as demons (from the point of view of the segments and the observer). In the classical theory, some power opens them, and they autonomously transfer ions in a potential-accelerated operating mode, and then that power puts the cap back on the top of the channel. The segments are separated by a semipermeable membrane having capped ion channels inserted into its material. As long as the caps of the ion channels are closed and the ion concentration on the two sides of the membrane is the same, we do not see any change in the state of the solutions. Although the channels can stochastically open, close, and reopen, they transmit more or less well-defined charge quanta. Even the channels can recognize the ions’ chemical nature and transmit only a selected ion type. The channels are passive during those processes, although the enormous voltage gradient can rearrange their structure and change their behavior through that. The demons also coordinate their actions using the layer containing charges as a communication medium; their population maintains a well-defined macroscopic current across the membrane.
2.11.2 Passing through ion channels
In Fig. 4, an ion channel is depicted in the middle of the figure with a diameter of about 3nm3\ nm for visibility. Furthermore, we assume that the ion’s size and, correspondingly, the thickness of the atomic layer in the electrolyte on the surface of the membrane is approximately 0.1nm0.1\ nm (although with their electric double layers, they can also reach a size of 1nm1\ nm). For comparison, recall that an Atomic Force Microscope (AFM) tip’s size is typically about 10nm10\ nm, the size of the tip of the clamp pipette is in the range of 1,000nm1,000\ nm and the size of the soma in the range of 10,000nm10,000\ nm.
Figure 4 also hints at how ions can move in the proximity of the membrane. On the left side, the electric field increases toward the membrane, so the ions cannot move in that direction without an initial force being applied. The case is the same on the right side because of the opposite charge. An ion must gain energy to get closer to the membrane, which means, at the same time, higher potential energy. This situation explains why the ions do not diffuse into the bulk region. At their present positions, the counterforces by the membrane balance the attraction from the opposite side. The ion channels can work in continuous mode (the "resting ion channels") and impulse mode (the "transient ion channels" are gated by "caps"): "Two types of ion channels—resting and gated— have distinctive roles in neuronal signaling. Resting channels are primarily important in maintaining the resting membrane potential, the electrical potential across the membrane in the absence of signaling. Some types of resting channels are constitutively open and are not gated by changes in membrane voltage; other types are gated by voltage but can open at the negative resting potential of neurons. Most voltage-gated channels, in contrast, are closed when the membrane is at rest and require membrane depolarization to open.[1], page 126.
Notice that the resulting electric field cannot accelerate an ion except when the ion is at the beginning of the open ion channel across the membrane. Here, the mechanical counterforce is missing, and the field drops to the value generated by the condenser inside the plates. Here, a vast electric field accelerates the ions to a high speed (although the force needed to move against the viscosity decreases the electrostatic acceleration). The empty red circles show an ion that traverses through the channel. The figure shows the electric fields on both sides of the channel at the beginning and end. (On the right side, the electric field is negative, but so is the charge of the ions.)
The electric acceleration sharply decreases after leaving the ion channel (see the right side). As shown by the difference in the vertical positions of the filled red and green balls, the difference in the electric field values is huge across the membrane ($E_{accelerate}$) and considerable inside the dielectric electrolyte ($E_{assist}$) as shown by the difference in the electric field values of the green balls. The field still accelerates, but significantly less intensely than the one in the gap. Suppose we assume that the ion quickly accelerates to its Stokes-Einstein speed. In that case, the travel speeds in those space regions are proportional to the difference of the fields in the shown positions and the reciprocals of their travel times.
The passed ions decrease the concentration and the electric field on the high-concentration side, and, conversely, increase them on the low-concentration side. At the exit, the field continues to accelerate the ion. However, the increasing potential and concentration increasingly decelerate the passing ions, so they quickly brake them to the assisted speed. Then, the effect of the electric field cancels, and the ions will move with their diffusion speed ("the ion stops"). The local potentials, connected to the operation of the ion channels, act as an "electric lock".
We note that moving charge inside the ion channels (nanopores) in membranes is a complex phenomenon and its details are discussed in [19]. They have studied theoretically and experimentally the generation of membrane potential at zero current in nanoporous membranes separating two reservoirs with different electrolyte concentrations. Although not specifically for our environment, they confirmed our statement that the membrane potential is of electrostatic origin and in the order of a couple of $mV$. Their finding that "the corresponding Donnan potentials appear at the pore entrance and exit, leading to a dramatic enhancement of membrane potential in comparison with an uncharged dielectric membrane" inspires beginning further research in uncovering quantitatively the details of the operation of the dynamic layer governing neuronal operation; furthermore, moving toward studying non-equilibrium processes.
We want to describe the change of voltage due to limited resources using an idea similar to the predator-prey model. We have charged sheets (represented by ions in the electrolyte) with potential $U_{H}$ and $U_{L}$ on the two surfaces of the membrane, plus we consider the corresponding neighboring layers on their side toward the "bulk" of the segment. We assume that the layers’ capacity is constant and has the same value, so the change in charge is linearly proportional to the change inf the voltage.
Figure 7: The potential values connected to the operation of ion channels. The arbitrary values of the transfer speeds (for visibility, instead of the real values) are $\alpha=1$; $\beta=.1$; $\gamma=0.0001$; furthermore the time scale is also arbitrary.
We consider that a charge transfer happens from the layer with potential $U_{H}$ to the layer with potential $U_{L}$  (that is, the same charge is removed and added, respectively) with a high speed (we call it potential-accelerated speed); furthermore, that with a much lower speed (we call it potential-assisted speed) the charge in the neighboring layer increases, and decreases, respectively, and those voltages are ($U_{H-1}$ and $U_{L+1}$).
That is, we assume four voltages and the initial conditions
$U_{H-1}=U_{H}=U;\quad U_{L=1}=U_{L+1}=0;$
We assume a constant capacity for the layers and that the amount of transferred charge (or voltage) is proportional to the difference in voltages in the respective layers.
$\frac{dU_{H}}{dt}=-\alpha*(U_{H}-U_{L})+\beta*(U_{H-1}-U_{H})$
$\frac{dU_{L}}{dt}=+\alpha*(U_{H}-U_{L})-\beta*(U_{L}-U_{L+1})$
Furthermore, we assume that the after-diffusion from and to the layers next to the proximal layers is negligible aside from the high speed of the charge exchange between the proximal layers, that is
$\frac{dU_{H-1}}{dt}=-\gamma*(U_{H-1}-U_{H})$
$\frac{dU_{L+1}}{dt}=+\gamma*(U_{L}-U_{L+1})$
As Fig. 7 depicts, voltages $U_{H}$ and $U_{L}$ quickly approach their balanced values. When they get equal, the driving force cancels. In the meantime, the voltages $U_{H-1}$ and $U_{L+1}$ tend to approach $U_{H}$ and $U_{L}$, respectively. The diffusional and energy-producing processes, depending on the local conditions, change the voltages on the two ends of the channel, creating the illusion that the channel opens and closes (ad-hoc constants are used to illustrate the concept). As [27] analyzed, this on/off behavior of currents can be observed for natural and synthetic lipid membranes, without caps. Despite slightly different experimental conditions, the amplitudes and typical time scales are similar.
2.11.3 Ion selectivity
The mechanism of channel passing can also contribute to explaining ion selectivity. "The normal selectivity cannot be explained by pore size, because $Na^{+}$ is smaller than $K^{+}$ [45]". The fact that ion selectivity changes with the membrane potential is observed (under the name of changed specific permeability), but it remains unexplained. "For example, the negative resting potential of nerve cells is largely determined by a class of $K^{+}$ channels that are 100-fold more permeable to $K^{+}$ than to $Na^{+}$. In contrast, during the action potential a class of $Na^{+}$ channels is activated that are 10- to 20-fold more permeable to $Na^{+}$ than to $K^{+}$. Thus a key to the great versatility of neuronal signaling is the regulated activation of different classes of ion channels, each of which is selective for specific ions." [1], page 101. However, the question of why the permeability changes and the channels are activated remains open. It is remarkable that an ion-specific channel is permeable also for other ions, only different transmission intensity is observed.
The commonly used picture of the operation of selectivity filters is surely wrong. The assumed mechanical operation of the pores is too slow: the assumed structural change needs $10^{-8}\ s$ and the ions passage time is about $10^{-11}\ s$ (furthermore, it must be repeated about $10^{3}$ times per passage). If a wrong ion is caught, it must be transported back to its departure side, through the right ions (against their repulsion), and the right ions must retry. Neither moving forward nor backward is an appropriate driving force present in that picture. Furthermore, the transport capacity is not sufficient for the operation.
In our model, as Eq.(6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamic forces are equal. The $E_{thermal}^{C_{k}}(d)$ "electric field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions are different. While a resulting force is exerted on one ion and moves it to the other segment, the other remains resting, or even moves in the opposite direction. Furthermore, the rush-in of ions greatly increases the $Na^{+}$ concentration in the dynamic layer, naturally increasing the number of $Na^{+}$ ions passing through the channel. That increase alone creates the illusion of the increased $Na^{+}$ permeability. Of course, the ions moving across the segments change the field contributions, indirectly affecting the balanced state of the other ions. The permanently open ion channels only passively participate in the process, and their apparent penetrability changes with the changing fields. To provide channel penetrability data is meaningful only if ion concentrations and membrane thickness data are also available. The driving forces of the ions are different, and they produce, depending on the actual concentrations and actual local potential, the selectivity of the channels.
2.11.4 Ion pumps
"These pumps differ from ion channels in two important details. First, whereas open ion channels have a continuous water-filled pathway through which ions flow unimpeded from one side of the membrane to the other, each time a pump moves an ion, or a group of a few ions, across the membrane, it must undergo a series of conformational changes. As a result, the rate of ion flow through pumps is 100 to 100,000 times slower than through channels. Second, pumps that maintain ion gradients use energy, often in the form of ATP, to transport ions against their electrical and chemical gradients. Such ion movements are termed active transport." [1], page 101. We add that, finally, the energy source of the ion channels is also ATP. However, the potential energy acquired by the condenser (in the form of creating resting potential) is prepared "globally" (not at the proximity where it is used), and during the full life cycle of neuronal operation, mainly in the recovery period.
The role of ion pumps must also be revisited. It is not clear what a force can "transport ions against their electrical and chemical gradients", and how can they deliver the ions in sufficiently short times needed for the operation. In the resting state, instead of some magic lipid mechanism (ion pumps), the resultant potential (the sum of the electric and thermodynamic potential for the individual ions) moves the ions through the channels in the membrane. The resulting conductance on those channels is about 50 times smaller [10, 12] than that of the AIS. In a typical mammalian cell, see Table 1, the sum of the electric and thermodynamical potentials is $32\ [mV]\ K^{+}$ ions and $4\ [mV]\ Na^{+}$ ions. The same values for squids are $33\ [mV]\ K^{+}$ and $13\ [mV]\ Na^{+}$. Correspondingly, in the resting state, the resulting force moves sodium ions out of the cell and potassium ions into the cell (for more numeric examples see section 4.1). The so called "Na-K pump" is supposed to actively transport (using energy from ATP) sodium ions out of the cell and potassium ions into the cell. However, no mechanism is provided for how the ion pumps gain energy from ATP and forward it.
As Figure 6 in [15] displays, the ratio of $K+$ and $Na+$ changes sharply during the transient state. Theoretically, they assume that the neuron “pumps 3 $Na^{+}$ ions out of the cell and two potassium ions in”; experimentally, they show in their Fig. 6 that the ratio changes between 0.01 and 7.5. According to our theory, the magnitude of the current (and so the exchange ratio) depends on the resulting potentials, so the ratio must change as the membrane potential changes during AP. The classical theory cannot explain this behavior and leads to exciting conclusions: "Furthermore, we analyzed energy properties of each ion channel and found that, under the two circumstances, power synchronization of ion channels and energy utilization ratio have significant differences. This is particularly true of the energy utilization ratio, which can rise to above 100% during subthreshold activity." [15]
Our model says that both the concentration (due to $Na^{+}$ rush-in) and local membrane potential drastically change during the transient state. As long as the resulting potential is above $\approx 33\ mV$, the driving force acting on potassium ions is positive, i.e., the pump will move both ions out of the cell. The potential gradient near the membrane also explains the energy delivery: ATP, by hydrolysis, generates ions which move to the "plates" of the condenser. (as we discuss in section 5, in the transient state, the "setpoint" of the control circuit is also changed, and to restore the condenser’s voltage, those ions must be collected.)
2.11.5 Delivering current across the membrane
The passage is too quick to affect the bulk, given that the ions can only use a potential-assisted speed to reach distant places in both segments. Again, the charge and mass conservation works: the ions pass suddenly from the high-concentration side to the low-concentration side, only from one layer to another. The mentioned layers on the two sides will actively initiate and terminate the ion transfer through the ion channels, but the ions can only pass through an open channel. One layer saturates, and the other empties. After a while, the source of ions will be exhausted. Those layers’ existence suggests revisiting the idea of describing neuronal operation by two single potentials of the bulks on the two sides of the membrane. The picture behind the equivalent circuits is definitely wrong.
Following their arrival, the driving force perpendicular to the membrane’s voltage disappears, and the ions form a thin "hot spot" in the layer. The electric repulsion acts in the direction parallel with the membrane’s surface and leads to distributing the ions (decreasing the gradient by distributing the charge locally) around the channel’s exit. However, the ions have finite speed. The ions saturate the layer on the membrane’s surface with a time constant between ($\frac{10^{-8}m}{10^{-1}m/s}\ s$) at the beginning and $(\frac{10^{-8}m}{10^{-4}m/s}\ s)$ at the end of their arrival period (we assumed 10nm10\ nm average distance between ion channel exits on the membrane). We shall take the longer time, so that we can expect a time constant for the saturation current around the ion channel’s exit in the order of 0.1ms0.1\ ms. When charging up the membrane in an avalanche-like way, the ions must pass on average a distance of about 0.05mm0.05\ $m$ from its center to its farthest point, so we expect a 0.5ms0.5\ ms ($\frac{5*10^{-5}m}{10^{-1}m/s}\ s$) time until the membrane’s slow current charges up the membrane to its maximum potential. The created charge must flow out from the farthest point in the neuron membrane of size 0.1mm0.1\ $m$ in time of order at or below 1ms1\ ms ($\frac{10^{-4}m}{10^{-1}m/s}\ s$); see the length of the $\frac{dV}{dt}$ pulse measured at the beginning of the AIS [54], which time is prolonged up to 10ms10\ ms by the neuronal $RC$ circuit; the ions are slow when the voltage on the AIS is low, see Eq. (25). Assuming those distances and speeds, including the potential-assisted speed of the slow current, we are on a time scale matching the available observations.
2.11.6 Synaptic gating
The slow current also explains the phenomenon of synaptic gating. The synaptic charge-up current flows into the membrane. It causes transient changes [55, 56] in the membrane’s voltage, providing direct evidence that the membrane is not always equipotential. The ions on the membrane’s surface can propagate at a finite speed. Given that the ions in the axonal arbor need to enter the membrane against the actual membrane potential, the membrane potential acts as a valve. The potential stops the ion inflow to the membrane for the period while the membrane’s voltage is above the threshold: it effectively inhibits further inflow through all synapses. This behavior naturally explains the absolute refractory period. After the membrane’s voltage drops below the threshold value, the ions can enter the membrane again (and the current intensity increases as the membrane’s potential decreases), but they need time to reach the AIS later when in the meantime the membrane’ voltage proceeded toward its hyperpolarized state; so they seem to appear dozens of microseconds later at the AIS, explaining the relative refractory period (as a delayed expansion of the absolute refractory period, with variable current intensity). Of course, the length of the relative refractory period sensitively depends on the conduction velocity [51].
2.11.7 Voltage sensing
"Voltage sensing by ion channels is the key event enabling the generation and propagation of electrical activity in excitable cells."[57] The mechanism by which voltage-gating of channels works remains a mystery; one of the worst consequences of the HH theory is that it separates the potential from ions and their current. It is not easy to investigate it experimentally: "the structural basis of voltage gating is uncertain because the resting state exists only at deeply negative membrane potentials" [58]. Usually, a "sliding helix" (structural) model is assumed.
Under certain conditions, an ion channel can be opened in only one direction and only for a limited period, and this allows the membrane to become semipermeable. We imagine an ion channel as a simple hole (a cylinder) between the high and low-concentration segments with a cap on its top (on the side of the low-concentration segment). At the spots where the ion channels are located, the ions cannot penetrate the membrane until the cap is removed/lifted (the channel gets open). Unlike the original Maxwell demon, our demon does not have information in advance about which particle should be transmitted: it is passive in selecting the particle. (Passive here means that no biologically produced local energy is used: the electric potential energy from the voltage difference across the membrane moves the ions to the other side of the membrane.)
Our voltage-controlled ion channel gets opened and closed due to purely electrostatic reasons. It works as a two-plate simple nano-scale electrometer (of type quadrant, Lindermann, Hoffman, and Wulf), similar to the ones used to measure the small electrical potential between charged elements (e.g., plates or fine quartz fibers). Given that the membrane and the cap in their resting state are isolators, no electrical repulsion is evoked between them, and the adhesion sticks them firmly together, representing a permanent force. The van der Waals force is inversely proportional to the squared distance between the dipoles in the cap and the membrane, respectively, and is linearly proportional to the perimeter of the channel.
However, when a slow ion current flows into the surface layer in the proximity of the cap, charges appear in the layer proximal to the membrane; the membrane and the cap get covered by a fragile electrical skin. The charge on the cap is proportional to the surface of the cap and similarly inversely proportional to the squared distance between the cap and the membrane. A local voltage gradient is generated by the local gradient of the slow ion current, and the force acting on the cap is proportional to the product of the voltage gradient and the area of the cap. Given that the cap is slightly elevated, the repulsion force may have a component in the direction of lifting the cap. Since the van der Waals force has a fixed size, when the electrical repulsion exceeds it at a critical voltage gradient value, the channel opens.
The gate remains open as long as the local charge distribution enables it. The cap is connected to the membrane only at one point, so it cannot fly away, and also cannot close again until the charge on the surface is present. In the absence of charge, the cap undergoes random movement, and the short-distance van der Waals force may eventually reposition the cap on the membrane, thereby closing the channel. The voltage-sensing electrometer opens the channel, and the lack of charge on the surface enables it to close, but the closing is not immediate (the mass of the cap is by orders of magnitude larger than the mass of an ion that can pass the channel). The fluctuation of the voltage gradient due to the slow current in the layer in the proximity of the membrane near the ion channel’s exit opens, closes, and re-opens the channel in an apparently stochastic way (actually, as the repulsion of charges due to the fluctuating current on the cap and the membrane regulates it), as observed.
When one cap is removed, the rushed-in ions in the proximity of the channel’s exit suddenly increase the local potential (produce fast transient changes [56]) proximal to the spot centered at the exit in the layer on the membrane’s surface. The surface outside the spot remains at a lower potential, so the ions in the layer start moving toward other channel exits, delivering potential to those channel exits. Given that they are voltage-controlled, they get open, and the process continues in an avalanche-like way [59]. The avalanche, as explained, needs a sufficiently large voltage gradient, which several synaptic inputs can trigger if they sum up appropriately. Alternatively, a single spike with sufficiently steep front slope [60] can be sufficient; providing a simple way of synchronizing neuronal assemblies (and proving that not the voltage, but the voltage gradient, single of summed, controls the operation).
The operation of the ion channel alone cannot explain why the channel closes after a given number of ions have passed the channel; that number is not (entirely) random. The local behavior of the membrane’s surface layers regulates the number of ions that pass through. The segments are no longer mechanically separated when the cap is removed. The charged ions are enabled to rush into the lower concentration segment. They experience an enormous accelerating gradient; that, comparable to that of electrostatic particle accelerators, "snorts" the ions from the high-concentration side into the low-concentration side and causes a process "like a flee hopping in a breeze" [25]. Consequently, "transport efficiency of ion channels is $10^{5}$ times greater than the fastest rate of transport mediated by any known carrier protein" [45]. Having gates (caps) is needed only for the synchronized operation of the channels: opening and closing work also in the absence of caps, as discussed by [27].
The snorted ions "hop" into the layer from another layer. In the beginning, with their voltage-accelerated speed, it could take less than $\frac{5*10^{-9}m}{10^{3}m/s}\ s$ to pass the channel (simulation [61] uses a psecpsec representative time interval), in the end, they may slow down to the voltage-assisted level as the potential gradually decreases (which is still $\frac{5*10^{-9}m}{10^{-1}m/s}\ s$), so we can omit that time when calculating the charged layer formation (as discussed in Appendix A, the estimated speed is about $10^{5}\ m/s$). Due to the enormous speed difference between the accelerated and assisted speeds, the passage is practically instant. On the high-concentration segment, only the ions in the layer in the immediate proximity of the entrance can feel the accelerating potential and move with the potential-accelerated speed. The after-diffusion with the potential-assisted speed from the next neighboring layer in the high potential segment is by orders of magnitude slower than the passage through the hole with the potential-accelerated speed.
The accelerating potential around the channel’s exit gradually (but quickly) disappears (see Fig. 7) when the particle exits the ion channel (see the green ion in the figure), and the ion arrives at the bulk potential. It practically stops: it can continue only with its potential-assisted (later with drift) speed, which is several orders of magnitude lower. However, the rest of the ions in the channel are still accelerated through the channel, and somewhat later, they also land in the formerly low-concentration layer, further increasing its potential and concentration. The passed-through ions increase the local potential in the layer in the low-concentration segment and decrease the local potential in the layer in the high-concentration segment. Given that the after-diffusion speeds in the layers are limited, "as ion concentrations are increased, the flux of ions through a channel increases proportionally but then levels off (saturates) at a maximum rate" [45].
Here, the effect of the finite resources explicitly appears: about $10^{3}$ ions are transferred per channel. These ions are snorted from one layer in the high-concentration segment into another layer in the low-concentration layer. The driving force gradually decreases because ions leave the first layer and appear in the second layer. The potential-assisted speed to replace the leaving ions in the first segment from the bulk as well as diffusing out from the second segment without appropriate driving forces, this process is by orders of magnitude slower; therefore, we can approximate the process as one where a gradually decreasing accelerating force drives the ions. In a very short period, the potential difference between the layers disappears (see Fig. 7) that prevents further ions from entering the formerly high-concentration layer: the process of transferring ions through the channel closes the door behind the needed amount of ions. Now the gradient diminished, and the van der Waals force can close the channel again.
The commonly used picture about the operation of ion channels [62] is definitely wrong, because
  • the potential generated across the membrane is entirely neglected
  • the ions have no driving force and the hypothesized protein carriers are too slow [45]
  • the considered van der Waals force is too weak to be noticed by the ions
  • the assumed force by the ’cation-attractive negative ends’ destabilizes the ion path: as the deviation from the central path increases, so increases the deviating driving force
  • even if the weak van der Waals force would work for a single ion, the next ion would be rejected by the strong Coulomb-force due to the first ion
3 Forming membrane’s potential
We model the neuron as an electric circle that comprises an electrolyte condenser implemented by its membranes, surrounded by two electrolyte segments. The membrane comprises low-transmission-capacity ion channels in its wall and high-transmission-capacity channels toward its axon; furthermore, it uses slow ions instead of electrons. The condenser produces a potential (that depends on the membrane’s thickness and the concentrations in the two segments, as discussed above). Furthermore, the two electrolyte segments produce two voltages, which are combined from Nernst voltages. According to Kirchhoff’s Voltage Law, in a balanced state, the sum of those voltages must be zero. However, one must take into account the effects of the "slow current" (emulated by "fast current" with a plausible time course). We omit the biological regenerative processes (such as ATP producing ions through hydrolysis, ion pumps and channels functioning, and so on) and discuss only the most important physical primary processes. In light of our findings, we must reconsider several key concepts in neurophysiology.
3.1 States
The two electric fields show significantly different dependencies on the neuron’s parameters. Recall that although the parameters we used are biologically plausible, they are only estimations. The qualitative conclusions, however, remain valid. Figure 1 shows, in the function of the membrane’s thickness, the electric field across the membrane due to electric charges (it does not depend on the thickness of the membrane), for different physically plausible assumptions; furthermore, the thermal electric field.
The balanced state is set where the thermodynamic and electrical "electrical field" diagram lines cross each other; around 5nm5\ nm; the value we assumed in evaluating our equations. On the other hand, we can check the electric field’s dependence on the summed-up concentration of the ions in the segment (see Fig. 8) (dehydration, for example, changes the setpoint through changing concentrations). Here, the thermodynamic electric field is constant but different at different width parameters. The field of the classic condenser does not provide a reasonably balanced state (a matching line). The dielectric diagram line seems to provide a realistic estimation when using biologically reasonable parameters. One can describe the neuron as a balanced system at the intersection of the sloping and the horizontal lines. In the resting state, only minor deviations (fluctuations) from this state exist in both the electric field and the chemical concentration.
Figure 8: The electric and the thermal "electric fields" at different concentrations, see Eq.(18). The thermal electric field varies with the thickness parameter and depends on the concentration; see Eq.(4). Balanced states are where the sloping lines cross the horizontal ones. The dot marks where $H$ provided measured data.
In the transient state, if we assume that a charge transfer causes a significant potential change around 100mV100\ mV in the layer of width 10nm10\ nm, it means a $10^{7}\bigl[\frac{V}{m}\bigr]$ jump in the electric field that corresponds to a $100\bigl[mM\bigr]$ jump in the concentration (for only a short time and only in that mentioned layer; but not in the bulk; see also Fig. 10). One can infer that the neuron can minimize its offset voltage (and so: its energy) by either decreasing its membrane’s width, the sum concentration of the electrolyte, the electric field, or a process with a temporal course that changes their combination, thereby implementing a true multi-physics phenomenon. During issuing an AP, the system finds its way back to the crossing point (the path is not simple: concentrations on both sides of the membrane and the electric field in the mentioned layers change, all having different speeds; electric, thermodynamical, mechanical and so on, changes happen; the process is a relaxation (a waving), as it is seen in the shape of an AP).
We can imagine how the electric process happens in this controlled system (for visibility, the positions of the marks are not proportional to the mentioned numbers). The system is in a balanced state at a crossing point. The two circles connected by bent arrows represent that when an excitation begins at the point with lower concentration and higher potential (the cytoplasm side and negative resting potential), ions rush into the surface layer, so the ion concentration increases, which increases the electric field; this is seen that the membrane potential suddenly increases. (We just note here, that, due to ions’ repulsion, the pressure also increases, proportionally with the potential.) The increased potential starts a current, but the current is slow, and the intensity of the current through the channels in the membrane’s wall is much lower than the current of the drain (AIS) so the system issues an AP while returning to its starting point. Given that it receives a large amount of $Na^{+}$ ions, its primary effort is directed to get rid of them (the electric gradient cannot distinguish the electric charges by their chemical nature), so the output current through the AIS fundamentally comprises $Na^{+}$ ions. The system will move along the line, with the restriction that it is a movement in the phase space (the change in concentration is not proportional to time), and the different points on the membrane’s surface may have the potential at different times (the spatiotemporal behavior). Due to the capacitive current, the local potential may temporarily drop below the set-point (as observed in physiology, this phenomenon is known as hyperpolarization). The time course of the processes is described by the laws of motion of biology [8] (the time derivatives of the Nernst-Planck equation). The simplified model operates as described in section 4 of [63] and its computational details are provided in [64].
3.1.1 Resting state
In the resting state, the goal is to provide stability by using slow and less intense currents. As [1] discusses, there are ’holes’ (non-controlled ion channels) in the membrane where the counterforce is missing. The difference in the electric and thermodynamic forces could accelerate ions until the Stokes-Einstein speed (see Eq.(27)) is reached. Given that the deviation from the set-point is slight, the intensity and speed of the current are low. The slow ion transport delivers both charge and mass, so the gradients permanently direct the process control variable toward the reference point. Given that the ’resting channels’ are scattered on the surface of the membrane and it takes time for the gradients delivered by slow ions to reach the entrance of an ion channel, some slight variations in the actual value of the membrane’s potential necessarily exist. Given that the AIS comprises non-gated ion channels with much higher density than those in the membrane, the variations and the control ("leaking") are marginal, and mainly through the ion channels in the membrane’s walls, as observed in experimental physiology.
The electric voltage contribution is unidirectional; the two-component thermodynamic contributions, depending on the concentrations, can be positive or negative. If choosing the zero potential at the middle of the electric potential, we can interpret that the concentration combinations maintain their ionic concentrations (and so also thermodynamic voltage contributions) independently on both sides, starting from a zero potential level. They generate potentials autonomously in the segments proximal to the opposite surfaces of the membrane. The electric charge on the membrane creates another voltage independently. However, that voltage must fit between the two thermodynamic voltages. Any deviation in the potentials starts a current, directing the concentrations and voltages toward a balanced state. The process is complex: the sum of surface concentration of ions must be the same on both sides for the condenser, and the ratios on the intracellular and extracellular sides must be adjusted to satisfy the conditions
$U_{internal}^{K^{+},Na^{+}}$ $=-U_{electric}$ $+U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}}$ $(+U_{internal}^{invasion})$ (33)
$U_{external}^{Cl^{-},Ca^{2+}}$ $=U_{electric}$ $-U_{external}^{Cl^{-}}$ $-U_{external}^{Ca^{2+}}$ $(+U_{external}^{invasion})$ (34)
In a balanced state, the left sides are zero: at the contact point between the electrolyte and the membrane, there are no potential differences, while in a perturbed state, they are non-zero and so they represent the driving forces for restoring the balance. (Recall that the ions are slow and that the local electric and concentration gradients control the local potentials.) From the derivation, it follows that
$U_{electric}$ $=-2*(U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}})$ (35)
$U_{electric}$ $=-2*(U_{external}^{Cl^{-}}$ $+U_{external}^{Ca^{2+}})$ (36)
$U_{electric}$ $=U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}}$ $=-(U_{external}^{Cl^{-}}$ $+U_{external}^{Ca^{2+}})\quad$ (37)
Our statement is the general formulation of that "Thus ions are subject to two forces driving them across the membrane: (1) a chemical driving force, a function of the concentration gradient across the membrane, and (2) an electrical driving force, a function of the electrical potential difference across the membrane." [1], page 129 and that "nerve and muscle cells have a particularly rich variety and high density of membrane ion channels, their channels do not differ fundamentally", page 107. Fig. 9 quantitatively underpins that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[65]. In all species, the chemical and thermodynamic processes keep balance in the cells of different construction in both resting and transient states.
According to the above derivation, we consider three characteristic segments in the neuron. In the middle, we have an electrostatically charged isolator plate that has positive and negative charge carriers on its two surfaces, and no charge carriers inside, so it has a homogeneous electric field inside it. The magnitude of the field depends only on the total concentration of the corresponding ions in the segments. On its two sides, there are ionic solutions of different compositions. If the membrane is permeable for the ions, they can move from one segment to the other as the resultant field dictates, until the fields get balanced.
Figure 9: The mechanism for controlling neuronal operation is the same in all cases shown in Table 1. It appears that the thicker membrane produces a higher membrane electric potential and requires a higher thermal potential to compensate for it. However, it can solve the task using a lower salt concentration, given that the electric membrane potential accelerates the ions across the ion channels in the membrane, and the operation is faster. The electric fields for the different species have similar values, as shown in Fig. 1. Furthermore, the total concentration decreases as the membrane thickness and potential increase.
The membrane’s electric potential defines the resting potential, and the "electric" and "thermodynamic" potentials work together to control neuronal operation. The anatomically defined "electric" membrane potential, plus the vast amount of ions in the extracellular segment, provides a stable concentration that serves as a solid reference point. In the resting state, the membrane potential is balanced by the presence of always-open "resting ion channels". When ions flow into the membrane, the "resting ion channels" can keep the balance: the internal $K^{+}$ and $Na^{+}$ concentrations are adjusted as the $Na^{+}$ flows in.
Figure 9 also shows that in its native state, the neuron is electrically balanced: in the internal and external segments, the resulting thermodynamical and electrical forces are equal. Any foreign invasion into the neuron’s life (chemical, mechanical, or electrical, including clamping) changes the left-side reference value to an externally set value. The neuron must adjust its $Na^{+}$ and $K^{+}$ concentrations to reach a new balance while the invasion persists. That is, any invasion ("testing in the physical laboratory") changes the concentrations and voltages inside the cell. In addition, the internal biological processes can trigger internal processes that contribute (from the measurement’s point of view) foreign currents or processes, see also section 3.1.2. The feedback from the clamping methods introduces a compensating current. Even a simple conductance measurement device is active: it applies a test voltage that generates a test current, which sensitively affects the ionic composition of the neuron (recall that the concentrations may differ by up to six orders os magnitude). Although modern electronic devices attempt to conceal this effect, they subtly influence biological operation.
From our result, it follows that for more chemical elements, a per-element coupled set of equations exists, plus the sum concentration agreed electric fields that are valid simultaneously and define the process variable (aka membrane potential) of the neural regulatory circuit. In a more symmetrical form
$0$ $=\frac{U_{electric}^{Total}(c,\Delta z,d)}{2}+\sum_{k}U_{thermal}^{C_{k}}(d)+{invasion}\quad for\ all\ k;\ inside\ and\ outside$ (38)
$c$ $=\sum_{k}C_{k}^{ext}=\sum_{k}C_{k}^{int}$ (39)
That is, we have an unidirectional electric potential difference and oppositely directed thermal voltages. In the balanced state, the two potentials must be equal. When changing say one concentration, it changes the other concentration on the same side, and the surface charge density of the condenser. The changed electric voltage re-adjusts the thermodynamic voltage, and as a consequence, the concentrations, in the other segment. There is no simple way to express, say, $Na^{+}]_{i}$ in the function of $K^{+}]_{i}$ or vice versa, and similarly for the other ions. Everything moves. The size matters: the amount of ions in the external segment is many orders of magnitude higher, and the gradients are reciprocally proportional to the amount of ions.
As the figure suggests, an external invasion, typically, an electric voltage on the intracellular side, changes the balanced state, and due to the parameters being linked, changes all concentrations. When moving the system out of its balanced state, in any way, a driving force appears that moves the system towards finding a new balanced state. However, recall that the ions are slow, so the changes are not instant.
Table 1 shows our results for three different famous published cases (Table 2.1 in[30]). It looks like the calibration of the concentration of negative and positive ions needs scaling, and the measurement of $Ca^{2+}$ comprises issues (the $Ca^{2+}$ concentration in all cases is inaccurate and had to be corrected).
Table 1: The summary data table for equilibrium (Data from Table 2.1 in[30] )
Ion Inside Outside $U_{therm}\frac{RT}{zF}log\bigl(\frac{[C]_{o}}{[C]_{i}}\bigr)$ $U_{electr}^{theor}$ $U_{electr}^{exper}$ $U_{membr}$ Width $E_{membr}$
[mM] [mM] [mV] [mV] [mV] [mV] [nm] [MV/m]
Squid (HH)  T=293 $K^{o}$ 3 14.0
$K^{+}$ 400 20 $75.5=58log\bigl(\frac{20}{400}\bigr)$ 41.9 41.4 -62.0
$Na^{+}$ 50 440 $54.8=58log\bigl(\frac{440}{50}\bigr)$
$Cl^{-}$ 40 560 $66.5=58log\bigl(\frac{560}{40}\bigr)$ 41.9 42.0 63.0
$Ca^{2+}$ $0.27^{*}$ 10 $45.5=29log\bigl(\frac{10}{0.27}\bigr)$
used for the published value 0.4
Frog muscle (Conway)  T=293 $K^{o}$ 8 9.61
$K^{+}$ 124 2.25 $101=58log\bigl(\frac{2.25}{124}\bigr)$ 88.9 83.6 -125.4
$Na^{+}$ 10.4 109 $59.2=58log\bigl(\frac{109}{10.4}\bigr)$
$Cl^{-}$ 1.5 77.5 $99.4=-58log\bigl(\frac{77.5}{1.5}\bigr)$ 88.9 82.7 124.1
$Ca^{2+}$ $0.021^{*}$ 2.1 $58.0=29log\bigl(\frac{2.5}{0.25}\bigr)$
used for the published value 4.9
Typical mammalian cell  T=310 $K^{o}$ 6 11.1
$K^{+}$ 140 5 $89.7=62log\bigl(\frac{5}{140}\bigr)$ 57.7 57.3 -85.8
$Na^{+}$ 15 145 $61.1=62log\bigl(\frac{145}{15}\bigr)$
$Cl^{-}$ 4 110 $89.2=-62log\bigl(\frac{110}{4}\bigr)$ 57.7 56.9 85.9
$Ca^{2+}$ $0.04^{*}$ 5 $60.8=31log\bigl(\frac{5}{0.04}\bigr)$
used for the published value 0.0001
$Na^{+}$ ions only [5] T=310 $K^{o}$ 6 10.0
$Na^{+}$ 12 145 $66.6=62log\bigl(\frac{145}{12}\bigr)$ 60.7 66.6
When calculating the values in Table 1, we used $\Delta z=1.05$ and employed the membrane’s thickness data from various publications to demonstrate that our theoretical approach is correct and that using the correct data can provide even the absolute values of the membrane potential. The higher-than-expected value of $\Delta z$ may suggest that the ions and membrane’s double layers in the electrolyte may also play a role. Dedicated, complete investigations can reveal the details.
3.1.2 Transient state
Figure 10: The mechanism of producing AP (using numbers referring to the case of squid shown in Table 1). In the resting state, the inside thermal potentials follow the green paths. When the rush-in of $Na^{+}$ ions suddenly increases the internal concentration to $[200]$, the potential on the internal side of the membrane increases to +90mV+90\ mV. The neuron must decrease its concentrations to restore the resting potential, mainly by releasing an action potential through the AIS.
In a special type of invasion, when suddenly a large amount of $Na^{+}$ rushes in (i.e., the trigger is external, but the ions derive from a biological process), the current throughput of the "resting ion channels" is not sufficient anymore: the internal $Na^{+}$ concentration in the proximal layer drastically increases, and so does the membrane’s potential. At that point, the neuron is out of balance: although most of the $Na^{+}$ ions can flow out through the AIS, the system also changes its $K^{+}$ concentration through the limited capacity "resting ion channels" until the balanced state is regained. This latter effect must not be mismatched with the capacitive current, which (due to its changed direction) causes the phenomenon known as "hyperpolarization".
Compare the figure to Fig. 7.2 in [1]. Notice the two-sided negative feedback effect: a change in the value of the "command potential" or "command concentration" changes the other value. In the case of voltage invasion, the ratio of the ion concentrations also changes. In this case, the feedback amplifier is represented by the delicate balance of the electrical and thermodynamic force fields. In the case of clamping, another (electrical) feedback amplifier is introduced into the system, and they work against each other. The neuron attempts to restore its balance without external potential (it cannot distinguish the foreign potential from its membrane potential), and changes its concentrations accordingly. A vital difference is the speed of the feedback amplifier. The electrical one works with electrical speed, while the thermodynamic one with several orders lower thermodynamic speed. When switching (or changing) the external ("command") voltage, the neuron must first get into its "steady state". It takes time. "The command potential, which is selected by the experimenter and can be of any desired amplitude and waveform" [1]. However, if the measured data readings are not from a steady state, the measurement is wrong: the different amplifier speeds (different current speeds) surely distort the measured value.
Figure 10 shows the schematic course of potentials and concentrations during issuing an AP. Follow the green path: the concentration ratio of $Na^{+}$ and $K^{+}$ ions define a thermodynamic contribution in the intracellular segment. Follow the blue path: the $Cl^{-}$ and $Ca^{2+}$ ions define a thermodynamic contribution in the extracellular segment.
On the left side, two positive ions are in the solution, and the membrane potential (due to charge separation) rises toward the outside segment; the ions cannot go "uphill". Similarly, the $Cl^{-}$ ions cannot go uphill (due to their negative charge, the same potential is also rising for them). This single membrane potential keeps all ions in their segments. However, the $Ca^{2+}$ ions can travel "downhill". Calcium pumps needed: although in minimal concentration, $Ca^{2+}$ ions flow into the intracellular segment, and they must be pumped out (BTW: this is likely also the reason why measuring $Ca^{2+}$ concentration is inaccurate).
As we discussed above, the concentrations and voltages in the extracellular segment, with its vast amount of ions, remain unchanged in resting and transient states. Follow the red path: when $Na^{+}$ ions rush into the intracellular segment. They increase the concentration from [50mM][50\ mM] to (say) [200mM][200\ mM] (from point (Rest) to point (AP)). This sudden change increases the thermal contribution of $Na^{+}$ potential to [20mV][-20\ mV], but because the total charge on the membrane significantly increases, the resulting voltage on the intracellular side of the membrane increases to [90mV][90\ mV]. The neuron could decrease its potential to the resting value if it could decrease the $K^{+}$ concentration to [100mM][100\ mM]. However, the ion-delivering capacity of the "resting ion channels" is not sufficient: they are sized to maintain the resting state. (Furthermore, the electrical contribution appears instantly, while the thermodynamic one, due to the finite speed of ions, only with some delay [8].) Follow the dashed orange path: the large amount of rush-in $Na^{+}$ ions drastically increases the surface concentration on the membrane, so the potential suddenly changes to [90mV][90\ mV]. Follow the dotted orange path: the electric field across the membrane suddenly changes its direction from positive to negative (see the slopes of the solid and dotted orange lines) so that the excess positive ions can move "downhill" out of the intracellular space through the high-conductance ion channel array (AIS) at the end of the neuron. Its internal end is at the high membrane potential, while the external end is at the low potential of the extracellular segment.
The setpoint on the extracellular side remains fixed. On the intracellular side, at the beginning of the AP, the actual voltage is above that of the setpoint on the extracellular side, so at the two ends of the AIS, a positive driving force moves the ions toward the downstream neuron. Given that positive $Na^{+}$ ions leave the intracellular segment, its potential continuously decreases. At some point, the voltage drops below the setpoint of the extracellular side, and the electric field across the membrane reverses. The driving force also decreases, and a low-intensity current restores the balanced state to the previous setpoint. When measuring at the AIS, one observes that the potential rises suddenly, then decreases below the set-point (hyperpolarizes the membrane), then a decreasing current (flowing in the opposite direction due to the changed slope of the resulting potential (the changed electrical field across the membrane), but comprising the originally rushed-in $Na^{+}$ ions) slowly restores the resting potential. Here comes the finite speed of ions again. The ions can follow the potential changes with a delay due to their finite speed, causing an apparent delay in the time course of the AP’s current.
Models in neuroscience (as reviewed in [66]) almost entirely ignore these aspects. In our physical model, we see that the measurable membrane potential and current change in the function of the ions’ speed, the concentration, and its time derivative; furthermore, all mentioned quantities depend on the effective potential. As Eq.(3) shows, the "thermodynamic electric field" increases as the temperature increases, and raises the setpoint. As Eq.(27) shows, the current decreases as the temperature increases, in line with the experimental evidence [50, 67]. The amplitude of the action potential is decreased, given that the setpoint has increased and its duration is reduced, given that the current decreases.
4 The Action Potential
As discussed, a smaller number of ion channels is distributed over the surface of the membrane, and the overwhelming majority of ion channels is concentrated in the AIS.
4.1 The physical process
In the resting state, we have a condenser that has some low-conductance resistors (the resting ion channels) and a very low intensity (nearly constant) distributed current flows, under the effect of a slightly varying (nearly constant) membrane voltage. Due to the membrane potential, a current through the always-open channels flows on the dendritic surface and the resting channels’ conductivity represents a drain with sufficient transmission. The resting current practically does not reach the AIS: the ion channels along the current’s path practically "shunt" the current. This is a static state that classic neuroscience assumes: some static current in and some static (leakage) current out; no significant gradient. In this state, the correct model is a condenser with a parallelly switched resistor; an integrator-type $RC$ circuit, see Table 2 (although when approaching the threshold voltage, the AIS plays some role. The smaller gradient changes, such as subthreshold excitations, produce "mini-AP waves" [15], providing direct experimental proof that in this state the parallel $RC$ circuit is not correct). Some perturbations must be counterbalanced, but their amount does not exceed a predefined threshold.
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, and the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave controversially during the transient period. The actual voltage gradient may temporarily direct the chemical gradients in opposite directions. Phenomena such as ’over-relaxation’, in this sense, similar to ’hyperpolarization’, may happen.
Our results align with the observation (see caption of 11.22 in [45]), that the significant processes occur in a thin layer of the electrolyte proximal to the membrane surface. The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a significant change in the membrane potential. Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions in that layer. However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or in a few atomic layers on top of each other, depending on the concentration.
Suppose we assume that at the beginning of an AP a large amount of $Na^{+}$ ions are transferred from the extracellular to the intracellular side. In that case, their concentrations change from $145/15$ to $45/115$ and the corresponding thermodynamic voltage contribution changes from +61[mV]+61\ [mV] to 25[mV]-25\ [mV] (i.e., altogether a 86[mV]86\ [mV] sudden increase in membrane potential). For the $Na^{+}$ ions, the resultant potential changes from +4[mV]+4\ [mV] to 83[mV]-83\ [mV]. That means when the AP begins, the "Na-K pump" stops (if the resulting potential provides the driving force for the exchange pump. The only way for the neuron to remove the excess $Na^{+}$ ions is to generate a current toward the AIS (where the other end of the ion channels remained at the extracellular potential), until the $Na^{+}$-specific driving force disappears. The $K^{+}$-specific driving force, changes from 90[mV]-90\ [mV] to 176[mV]-176\ [mV], i.e., the $K^{+}$ intake gets more intensive, misleading researchers into believing that it causes the observed hyperpolarization. However, as Fig. 3 in [15] demonstrates, it occurs instantly, not with a delay; furthermore, the "leakage current" and the stimulus current are negligible. The low extracellular $K^{+}$ concentration plus the low number of channels in the membrane’s wall do not enable a significant increase in the intracellular $K^{+}$. (Given that complex changes occur, including changes in the electric potential that also change the concentrations, different waves start; the statement is not strictly valid.)
As the current flows, the thermodynamic force decreases, and at some point, the resultant force changes its sign; so does the direction of the current. This is seen as the "condenser current" (and mis-identified as an intense $K^{+}$ current). The ions move slowly, and due to the enormous electrostatic repulsion, their "electrostatic fluid" must be continuous. A special "damped wave" is formed (with all secondary mechanical, optical, etc. effects; see section 4.4) and the condenser current reverses. Electricity interprets the phenomenon as "capacitive current" in the equivalent circuit, and thermodynamics interprets it as changed chemical concentration (or the negative amplitude of the damped elastic oscillation). Our discussion adds that the electric repulsion generates a pressure with a shock wave in the neuron’s closed volume.
Figure 11: A summary of producing neuronal AP as a 3-D figure. The thick solid 3-D line shows how the color-coded voltage gradient drives the membrane potential in the function of the time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. These 2-D plots are delivered by the physiology; the figure explains how those parameters are interrelated. The red diagram line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 12. It produces the blue diagram line AP, see also the bottom inset in Fig. 12. The interdependence of the gradient and the voltage it produces leads to Fig. 14, left side, that allows to calculate the thermodynamical characteristics, right side, of neuronal operation.
The neuronal AP generation can be entirely described by the 3-D line in Fig. 11. The three dashed diagram lines are drawn in three perpendicular planes; see the figure’s caption and the discussion below for the details.
4.2 The oscillator model
Despite the early warning that ’it was not possible to separate the change into resistance and capacity components[31], a commonly accepted truism was that neurons, in some sense, behave as electric oscillators. HH introduced the idea explicitly that the electrically equivalent circuit of a neuron is an $RC$ oscillator. We must also add the invention from about three decades later: "Neurons ensure the directional propagation of signals throughout the nervous system. The functional asymmetry of neurons is supported by cellular compartmentation: the cell body and dendrites (somatodendritic compartment) receive synaptic inputs, and the axon propagates the action potentials that trigger synaptic release toward target cells. Between the cell body and the axon sits a unique compartment called the axon initial segment. The AIS was first described 50 years ago [i.e., nearly two decades after HH published their study], and its molecular composition and organization have been progressively elucidated during the following decades. …. Recent years have also brought crucial insights into the functions of the AIS: how ion channels at its surface generate and shape the action potential." [12] We provide the physics and mathematics of how AIS shapes the action potential (or more precisely, we show what an important role it plays in forming AP).
Inventing AIS changed the viewpoint of neuroscience [12], a half century later, after setting up the commonly accepted model by HH. "The axon initial segment (AIS) is located at the proximal axon and is the site of action potential initiation. This reflects the high density of ion channels found at the AIS. … The summation of synaptic inputs gives rise to action potentials at the axon initial segment (AIS), a 20–60 $\mu m$ long domain located at the proximal axon/soma interface that has a high density of voltage-gated ion channels." As discussed in [13], see also their Figure 1, the structure of the Axon Initial Segment is known to the most minor details. As the illuminating investigations in 2008 [10] revealed, the AIS has very dense ion channels. That is, from an electrical point of view, those parallelized channels can be abstracted as a discrete conductance (or resistance) between the membrane and the axon. The membrane itself can be abstracted as a distributed condenser with no resistance. This model is in contrast with the viewpoint of biophysics, that the membrane plus AIS is considered a distributed element, where the capacitor and condenser cannot be separated. The current model is the result of projecting the physical processes in the resting state to the transient state. Notice the important point: "Neurons are also anatomically polarized, as they can be subdivided into a somatodendritic input domain and an axonal output domain" [9]; providing direct evidence that (unlike in HH’s model) the input and output currents (and voltage time derivatives) are independent.
Table 2: RC circuit types
The RC Integrator The RC differentiator
$V_{out}^{Integrator}=\frac{1}{RC}\int_{0}^{t}V_{in}{dt}$ $V_{out}^{Differentiator}=RC\frac{dV_{in}}{dt}$
Low Pass Filter High Pass Filter
Resting state Transient state
[Uncaptioned image] [Uncaptioned image]
(Mostly) resting state (Mostly) transient state
Ion channels in the membrane Ion channels in the AIS
In our model, the AIS gets part of the membrane, and this leads to crucial changes. (BTW, the name is slightly misleading: the AIS is part of the neuronal oscillator, and it forwards a traveling potential wave to the axon instead of belonging to it.) ’Although by definition a neuron must have an axon to assemble an AIS, the relationship between AIS assembly and axon specification in vivo has not been determined yet’ [9]. Anyhow, a neuron can be abstracted (in the disciplinary view of electricity) as comprising a resistor-like and a capacitor-like component. However, in the resting and the transient states, they are combined in different ways.
In the resting state, we can model neurons as a parallel $RC$ circuit (see the "integrator" in Table 2), where "random" currents flow in and out. Practically, the output voltage is constant (the resting potential): the ion channels and pumps are distributed over the membrane’s surface, and due to the low speed of currents, at low local voltage gradients, the AIS is not involved. Only the synaptic inputs cause a measurable gradient on the AIS. In this native state, there is no gradient; in this sense, it is similar to the case of clamped operation, where the gradient is (apparently) suppressed by the feedback. HH did not see any structural elements on the membrane, so logically, they modeled it as a distributed resistor and capacitor, which really has resemblance with a parallel $RC$ oscillator. However, they made a wrong choice of the circuit type, and their choice (likely due to inertia) was repeated in good textbooks such as ([30] Figure 3.1 or [25] Figure 1.1), and it is a commonly accepted fallacy even today [68].
In the transient state, the high-transmission ion channels in the AIS (high conductance $R$) replace the low-transmission ion channels (low conductance $R$) in the membrane’s wall, and the rush-in current flows out entirely on the AIS (for the model, in this state, we omit the low-transmission ion drain). Those two elements form a serial $RC$ circuit (see the "differentiator" in Table 2) that acts as a damped oscillator (see section 4.4) and produces a current pulse, known as AP. Neuroscience mistakenly claims it is a parallel $RC$ circuit; that statement is valid only in the resting state. However, the number and intensity of the ’resting’ ion channels in the membrane’s wall are insignificant for producing AP, so we use the approximation that only the ’transient ion channels’ in the AIS are active when generating an AP.
In electric view, the process can be modeled as follows: a sudden voltage gradient (see Figure 12 middle inset) appears on the membrane (as a discrete capacitor), and a current flows out (see Figure 12 bottom inset) through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [49] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see section 4.3, furthermore, [63], section 4; for its algorithmic specifics [64], for further details [16].
Figure 12: The physical processes describing the membrane’s operation. The rushed-in $Na^{+}$ ions instantly increase the charge on the membrane, then the membrane’s capacity discharges (produces an exponentially decaying voltage derivative). The ions are created at different positions on the membrane, so they need different times to reach the AIS, where the current produces a peaking voltage derivative. The resulting voltage derivative, the sum of the two derivatives (the AIS current is outward), drives the oscillator. Its integration produces the membrane potential. When the membrane potential crosses the voltage threshold value, it switches the synaptic currents off/on.
The fundamental difference between the parallel and serial $RC$ circuits is that the parallel circuit cannot produce output voltage with opposite sign. As we discussed, the capacitive current, which by definitionem changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have enough transmission capacity. Not knowing about the AIS leads to assuming that the resting ion channels work also as transient channels (in other words, assuming a parallel $RC$ circuit instead of the serial one, furthermore, that the output current flows through the membrane instead of the axon). Some $K^{+}$ current certainly exists; for the magnitude of currents during AP, see [15].
The shape of the output waveform depends on the pulse width ratio to the $RC$ time constant. When $RC$ is much larger than the pulse width, the output waveform resembles the input signal, even with a square wave input. (In the case of the neuronal oscillator, the shape of the front side of the spike is similar to the one derived for the rush-in current, while the back side is very prolonged.)
4.3 Mathematics of the ActionPotential
Although the implementation of a neuronal circuit works with ionic currents and has a charge transmission mechanism drastically different from the one based on free electrons used in metals, the operation can align (to some measure) with the instant currents of physical electronics (as we suggest, the effect of the slow current and finite size can be reasonably imitated with current generators generating a special current shape, see section 2.10). The input currents in a biological circuit are a large rush-in current through the membrane in an extremely short time, and much smaller gated synaptic currents arriving at different times through the synapses. Those currents generate voltage gradients, and those gradients direct the operation of the $RC$ oscillator (Notice that due to the multi-physical effects, the parameters $R$ and $C$ may change in the function of time and the actual voltage). As it is well known from the theory of electric circuits, the output voltage measured on the output resistor is as follows:
$V_{out}^{Differentiator}=RC\frac{dV_{in}}{dt}$ (40)
where the input voltage is as follows:
$\frac{d}{dt}V_{in}=\sum\frac{d}{dt}V_{IN}^{Component}-\frac{d}{dt}V_{OUT}^{AIS}$ (41)
that is, the (temporally gated) sum of the input gradient that the currents generate, plus the gradient of the output current through the AIS [13, 12]. The latter term can be described as follows:
$\frac{d}{dt}V_{OUT}^{AIS}={\frac{1}{C}}\frac{V_{internal}-V_{external}}{R}$ (42)
The input currents are slow (we typically imitate them by a current generator with a special form), see section 2.10. In the simplest case, the resulting voltage derivative comprises only the rush-in contribution described by Equation (32) and the AIS contribution described by Equation (42). The gating implements a dynamically changing temporal window and dynamically changing synaptic weights. The appearance of the first arriving synaptic gradient starts the “Computing” stage. The rush-in gradient starts the “Delivering” stage [64]. The charge collected in that temporal window is as follows:
$Q(t)=\int_{t_{0,i}}^{{t_{thr}}}dt\underbrace{\sum_{i}}_{{t_{0,i}\leq t\leq t_{thr}}}I_{syn,i}(t)$ (43)
The total charge is integrated in a mathematically entirely unusual way: this is NOT a simple "integrate and fire": the collected charge defines the upper bound of the integration time. Through adjusting the integration time window, the neuron selects which one(s) out of its upstream neurons can contribute to the result; the individual contributions are the result of a bargain between the sender and the receiver. The upstream neuron decides when the time window begins, but the neuron decides when a contributing current terminates. The beginning of the time window is set by the arrival of a spike from one of the upstream neurons, and it is closed by the neuron when the charge integration results in a voltage exceeding the threshold. The speculations [33] about the spatial length of the spike are correct, but its "payload length" is defined by the receiving neuron. The computation (a weighted summation) is performed on the fly, excluding the obsolete operands (the idea of Analog In-Memory Computing (AIMC) [69] almost precisely reproduces the charge summation except for timing). The only operation that a neuron can perform is that unique integration. Its result is the reciprocal of the weighted sum, and it is represented by the time when a spike is sent out (when that charge on the membrane’s fixed capacity leads to reaching the threshold). Notice that the time is defined on the neuron’s local time scale and is interpreted similarly by the downstream neurons: the data transmission time matters.
4.4 Damped oscillation
According to Newton’s 2nd Law, a body with mass $m$, under the effect of a "spring force" proportional to its position $x$, and a linear damping force proportional to its velocity $v$ moves according to the equation
$\underbrace{m*\times\frac{d^{2}\,x}{{dt^{2}}}}_{moving\ force}+\underbrace{c\times\frac{dx}{dt}}_{viscous\ damping}+\underbrace{k\times x}_{spring\ force}=0$ (44)
The equation directly corresponds to our Eq.(7), given that neither constraint force nor external force is present. Recall that according to the Stokes-Einstein relation (see Eq.(25)), the viscosity force is proportional to the speed of the body in the fluid. The damped oscillation has an analytic solution (see Fig. 13 for the corresponding graph)
$x=e^{-\gamma*t}*a*\cos{(\omega*t-\alpha)}$ (45)
Figure 13: The analytical solution of the motion equation of the underdamped oscillator for some damping parameters.
A considerable resemblance with the shape of an AP is evident, but there exist considerable differences from the ’net’ physical case. Among others, the mass $m$ changes continuously; furthermore, it is distributed in the volume of the neuron, so one cannot expect a perfect description from such a simple motion equation. However, it is perfect for demonstrating that such a system can produce the symptoms of relaxation and hyperpolarization that physiology observes. (The electrostatic interaction is the strongest and fastest interaction in nature, stronger than the other interactions that have a role in the process, so we have reasons to think that Eq. (40) provides the best description.)
The "moving force" has two major components. The sudden rush-in of $Na^{+}$ ions has (at least) a double effect. One effect is that they press the surface of the membrane that elastically increases its diameter [70]. Due to their mutual repulsion, the ions experience a strong electric force toward the membrane. It represents an impulse $J=F\,\Delta t$ (N  s), that enables estimating the time and energy of the change the rush-in causes and explains that the energy needed to generate an AP is suddenly produced electrically and stored temporarily as elastic energy (the energy is produced in the background, when the ATP produces ions by hydrolysis, as we described above); see Appendix A. At the same time, it offers a plausible way of unifying the electric and thermodynamical attempts to describe neuronal operation; furthermore, a pictorial understanding of why the elastic change produces opposite amplitudes in the size of the membrane and the voltage in its surface layer. Similarly, it explains that in the first phase of generating an AP the neuron invests (performs a work $Q_{H}$ against the membrane’s elastic force), while in the second phase the electic membrane performs work $Q_{L}$ on pressing out the ions physiology observed in an AP; for the details see section 4.5. These two changes are inseparable, and their combined effect can explain the observed consequences. The compression and expansion give way to explaining the observed mechanical, optical, thermal, etc. changes.
Given that, as we explained, the energy operating the neuron is temporarily stored in the form of elastic energy, it is not surprising that its time course is in resemblance to a damped elastic oscillation. We can consider that the neuron forms a damped oscillator (the Stokes-Einstein force represents a viscous fluid, the $Na^{+}$ represents a $\delta$-function like excitation) with time constant about $\zeta=0.35$.
Their other effect is that the appearance of the ions drastically increases the potential in the surface layer of the neuron membrane. During regular operation, a very intense current pulse is created: "A small flow of ions carries sufficient charge to cause a large change in the membrane potential. The ions that give rise to the membrane potential lie in a thin (<1nm<1\ nm) surface layer close to the membrane" [45] (see the dotted line in Fig. 4; notice that the horizontal scale on the figure spans a few $nm$). For this, a considerable number of ion channels are used. When the trigger signal is received, the gated ion channels in the membrane’s wall open quickly and only for a brief period. The rushed-in ions create a large offset potential in the layer proximate to the membrane’s surface (from an electrical point of view, a condenser) that drives an intense current through the always-open ion channel array AIS (from an electrical point of view, a resistor) at the beginning of the axon. The slow current on the membrane’s surface mimics a condenser: in the picture of classical electricity, the charges stay inside the distributed condenser while they travel on its surface. In classical electronics, the condenser is sizeless.
4.5 Thermodynamic aspects
As the middle inset in Fig. 12 shows, the sum of the rush-in and AIS gradients controls generating the AP. During excitation, $Q_{X}$ energy is fed into the neuron. The excitation increases the neuron’s energy by $Q_{H}$. When the membrane potential reaches its peak value, the gradient changes its sign and starts to decrease the potential. In the transient state, the size of the neuron (due to the increased electrostatic repulsion) slightly increases [70] and stores the energy mainly in the form of elastic energy. The resultant of the electric, elastic and thermodynamic forces moves the ions toward the AIS. Although part of the energy is dissipated while the ions exit through the AIS with their Stokes-Einstein speed, the rest of the energy (the elastic part) is recovered. The concentration and potential values change as discussed in connection with Fig. 10. The membrane voltage controls synapses as shown in the top inset in Fig. 12.
Refer to caption Refer to caption
Figure 14: Thermodynamic consequences of the processes shown in Figure 12 during the generation of an Action Potential. Left: Comparing the measured (broken blue) and theoretical (red dotted) voltage gradient vs voltage phase diagrams [54]. Right: Reinterpreting the phase loop in terms of thermodynamics: the “Carnot cycle” (theoretical thermodynamic cycle) of a neuron
Figure 15: The time course of the electric gradient [54] perfectly matches that of the mechanical and thermal changes [27]; however, classic theory cannot explain the connection between them [8]. The experimental $\frac{dV}{dt}$ uses an opposite sign convention.
The ‘phase loop’ (how the voltage gradient and membrane voltage correlate; a parametric curve that uses time $t$ as parameter) shown on the left side of Fig. 14 represents a strong test of the theoretical line shape: the values of the potential and its gradient are simultaneously compared to the experimentally measured ones. Comparing the experimental and theoretical shapes supports the theory that describes the mutual dependence of voltage and its gradient on each other. Furthermore, it also shows that omitting the gradient from the classic theory misled the experiment designers, and the measurement design was wrong: the time interval between recording data points was too long. Where the gradient is high and changes quickly (see Fig. 12, middle inset), the blue experimental diagram line is roughly broken. The changes in the values are too significant, causing the integration to distort the value and resulting in an inaccurate broken line. In the rush-in region, the theoretical line is even out of range of the figure. Where the gradient change is slow, the measurement points are sufficiently dense, and the theory perfectly describes the experimental data. The classical theories cannot interpret the phenomenon and do not produce any similar dependence; furthermore, they affect the experiment design in the wrong direction.
An exciting aspect is that the phase plot in Fig. 14, left side, shows a remarkable resemblance to a specialized thermodynamic Carnot cycle in Fig. 14, right side; and inspires us to interpret the parameters of the "neuronal Carnot machine" (the bent arcs serve only for illustration; they do not follow the path on the left side). As discussed in [27], a neuron emits and absorbs heat (energy) in the first and second phases, respectively, of an AP; see Fig. 15. The experience, alone, defies the classical HH mechanism. Laws describing currents of electrons must not be used directly to currents of ions. Neglecting the repelling force between the charge carriers of the ion current, which can be omitted in the case of electron current, leads to entirely wrong conclusions, from the wrong oscillator model to losing the connection between the electrical and mechanical changes, to losing the connection between inanimate and living nature. Our theoretical model, unlike all the other models, can explain that observation quantitatively.
We rescale the horizontal axis with the resistance $R$ of the AIS (actually, the horizontal axis is transformed to $I_{AIS}$); thus, the plot becomes a neuronal energy density diagram. Actually, it corresponds to a usual PVP-V diagram, and we expect that the temperature changes in the same analogy (we keep the convention of using the letter $V$ to denote both potential and volume). By integrating the parametric $t$ values of any two points on the diagram ‘circle’ line, we receive the work performed by the neuron.
We consider that the environment in the period ‘Excite’ invests energy $Q_{X}$ through the neuron’s synapses (increases the membrane’s voltage to the threshold) between [0,1] to open the valves. Opening the valves enables the neuron receiving in period ‘Rushin’ energy $Q_{R}$ by letting in $Na^{+}$ ions, accelerated by the field across the membrane. Suddenly, the ion concentration and the membrane potential increase to their maximum value. Given that the process is instant, the potential $V$ does not change (no current flows), so the integral [1,2] evaluates to zero (an ‘iso-voltage’ transition). Now the neuron has a high potential to prepare its (already encoded) message (the AP). As our model explains, in the first phase (from the threshold potential to the maximum value of the AP, where $\frac{dV}{dt}$ is positive), the neuronal power cycle produces a positive heat $Q_{H}$, given that $\frac{dV}{dt}$ is always positive. In the second phase (from the top value of AP to where the driving force $\frac{dV}{dt}$ becomes zero again), the neuronal power cycle produces a negative heat $Q_{L}$, given that $\frac{dV}{dt}$ is always negative; in line with the measurement results shown in Fig. 15 [27]. The integral [2,3] provides the “heating” $Q_{H}$, the energy invested. Similarly, the integral [3,6] provides the “cooling” $Q_{L}$, representing the energy recovered (assuming no artificial current). In the “Relax” phase, $\frac{dV}{dt}$ is (almost) zero, so the corresponding integral evaluates to zero (‘iso-gradient’ transition). In this phase, the neuron must restore its resting state: by using ATP, it creates new ions. The difference in the energies of $Q_{H}$ and $Q_{L}$ is the net energy used to forward an AP; while $Q_{X}$ is the energy used to perform the computation. The energy $Q_{X}$ is an approximate value: as shown in Fig. 12, at the time of re-opening the synaptic inputs, the membrane can be above or below the resting potential, which actually means a neuron-level memory. In this sense, adjacent neural spikes can borrow energy from each other (this way explaining why the first spikes in a burst behave differently). The energy $Q_{R}$ is produced by the ATP hydrolysis in parallel with the ’primary processes’ described here, and the produced ions are acquired on the condenser plates, thus restoring the resting potential and providing potential energy for the forthcoming spikes.
From Fig. 14, after recalibration (with $R_{AIS}=10^{7}$ Ohm), one can estimate (by approximating roughly the area under the arcs by triangles, in units of $10^{-7}[J]$), $Q_{H}=5.2,Q_{L}=1.8,W=3.4,Q_{X}=0.2$. That is, according to the "phase plot" [54], the neuron operates with an efficiency around 65%, and approximately 3% of the energy is derived from the upstream neurons. The values (without using data from dedicated measurements) agree excellently with the measured [15] values $W=2.5$ and 74%; furthermore, the $Q_{X}$:total ratio $1:35$, derived by measuring the direct energy consumption [14].
In our model, the $Na^{+}$ rush-in increases the potential, the concentration, the pressure (due to electric repulsion; so it changes the membrane’s capacity), and the volume (since the membrane is elastic), so a multi-dimensional Carnot diagram could describe the state accurately (although the invested energy can be safely estimated from the work the charge performs given that the energy is almost entirely stored as elastic potential energy due to the enormous elasticity modulus). Furthermore, all parameters have a time course. It is not possible to provide a more accurate estimation without a true multi-physics simulation. However, the approach is promising.
4.6 Heat emission/absorption
In the framework of the classic theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}*R$) the ion currents flow. The existence of "cooling", alone, undermines the credibility of the classic theory [38]. The issue here is that the dissipation was attributed to the "leakage current." As Hodgkin wrote: "Hill and his colleagues found [71] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." [72] "All authors came to similar conclusions: during the action potential, no significant net heat is produced. Transient heat releases are mostly reabsorbed in the second phase of the action potential. …The finding of a reversible heat release during the action potential of nerves is a striking and very fundamental fact. It is inconsistent with the HH model. The physics underlying the nervous impulse must rather be based on reversible processes". [73] Our Speculations in Appendix A on the energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. (Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life".)
The heat liberation and absorption are direct experimental evidence of the "slow current" and the mechanisms described above. The lack of leakage current (see section 2.6.4) not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP. Pressure changes accompany the electric process and naturally explain, quantitatively, the observed temperature changes. As our explanation shows, only a small fragment of energy is stored in the form of electric potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [73]. The energy is almost exclusively stored as reversible elastic energy. The conclusion in [35], that "brain cells communicate with mechanical pulses, not electric signals", although shocking for the first reading, and is only half of the truth, is essentially correct. More precisely, mechanical pulses (pressure waves) deliver the charge carriers of the electric signals, in close interaction with the repulsion of the delivered charged carriers. In Equation (7), we also must consider the mechanical force delivered by the pressure wave. Furthermore, the nervous pulse transmission over the axons must be revisited. The enormous pressure, the incompressible electric fluid, and the repulsion of the delivered particles suggest an alternative and more reasonable mechanism.
4.7 Neural entropy
The theoretical diagram line in Fig. 14, based on the presented correct physical model, might serve as a good starting point for interpreting entropy for neuronal operation, again in analogy with the Carnot cycle; although the charge of the ions (the macroscopic current), complicates the process. Furthermore, material convection happens in different ill-defined volume parts of the membrane in different directions. Initially, in the resting state, the ions are in a (relatively) disordered state. After the rush-in, in the dynamic layer, the ions have a highly ordered macroscopic speed (see Stokes-Einstein speed, the gradually changing gradients, and the finite speed) component toward the AIS. In the layer, the maximum ordered state is reached when the voltage drop on the AIS reaches its peak value (the current is slow). Then, the ions slow down, and at zero voltage gradient, they reach again a minimally ordered state. Then, due to the ‘capacitive current’, the current changes its direction, and the backward current represents an ordered state again. Perhaps the latter-mentioned negative contribution is the one referred to as ‘negative entropy’ when describing life through thermodynamic processes? We must be aware that in the background, energy-producing processes are going on, comprising ordered material transport. Considering the neuron, based only on the described primary processes, would be a mistake.
The above discussion might help in understanding the physical entropy in the neuron. However, the information entropy is different. As our discussion underpins, the "information" arrives at the neuronal inputs in the form of charge carriers, and the time course of the input currents, in cooperation with the neural environment as well as the neuron itself, carries the information in some form. The information is distributed in space and time [74, 75], so we can be sure that using the pulse count as a measure of neural information is wrong. We must distinguish the carrier of the information, the ion, and the information content that its temporal appearance carries. After having the correct physical model of operation, we have a chance to gain more knowledge about how the information is encoded in the spatially and temporally distributed observable signals; furthermore, we have better hopes to understand the neural information processing, including the brain’s entropy handling.
5 Controling neuronal stages
From control theory, it is known that the goal of a controlled system is to govern the application of system inputs to drive the system to a desired state while minimizing delay, overshoot, steady-state error, and ensuring control stability. The neuron implements a controller that monitors the controlled process variable (membrane voltage) and compares it with the reference or setpoint (the resting potential). Recall that the geometry, and the composition and concentration of electrolytes, define the setpoint. The difference between the process variable’s actual and desired values, known as the error signal, represents the actual offset potential. It is applied as feedback to generate a control action to bring the controlled process variable to the same value as the setpoint.
5.1 Neuron as a PID controller
Biology uses such a simple Proportional-Integral-Derivative (PID) controller which has a proportional, an integrated and a derivative component. In simple words, it means that the output of the system (the output voltage measurable on the AIS) depends on the input voltage (input current if the resistance is constant), the integrated input current(s) and their derivatives (provided that the membrane’s capacity is constant). The sum of those summands defines the resulting output voltage. The biological case is more complicated than the technical one because biology works with slow ions and so those components have their temporal dependence, furthermore, the summands have several constituents and the neurons have an autonomous processing of those current-related constituents. That is, unlike assumed in the classic physiology, the output voltage is not necessarily a simple function of the mentioned input variables. The classic models all omit the derivatives’ contribution (a direct consequence of using ’clamping’) except that of the capacitive current. Furthermore, the ’Integrate and Fire’-type models consider the integrated currents only in the ’chargeup’ (’Computing’) period of operation and they entirely omit that different physical processes are going on in different stages of operation. That is, even the best classic models can not describe neuronal operation completely and correctly.
The gradients are used to adjust the process variable by their positive and negative contributions (corresponding to rising and falling currents), and the different speeds of the thermodynamic and electrical interactions minimize the delay (i.e., provide the maximum speed of operations, vital for survival). The steady-state error is minimized by setting the process variable to the reference point using long-term stable parameters (geometry and overall concentration). The low-intensity current through the always-open resting ion channels provides dynamic stability in the steady state.
We considered that the neuron has a stable base state. On the one side, this resting state must be dynamically stabilized for little perturbations using as little energy as possible (and to provide a mechanism when the cell grows, divides, or ages). On the other side, it must be able to restore the state after rough perturbations as quickly as possible, causing short-time transients (when restoring the membrane’s potential after issuing a spike). Both in the resting state and in the transient state, the system attempts to return to its balanced state, although this action in the two states require different physical mechanisms.
The system aims to provide an intense output that informs the downstream neurons. This overshoot is implemented by a large number of voltage-gated ion channels (distinct from the non-gated ion channels used to maintain the resting potential) distributed in the neuron’s membrane. The overshoot current flows through those persistently open ion channels, concentrated in AIS (which has about two orders of magnitude higher channel density than the wall). The intense slow current produces a condenser-like behavior (capacitive current); the phenomena called "polarization" and "hyperpolarization" (not polarization; instead, a movement of completely separated charge in the surface layer of the electrolyte) of the membrane provide the necessary positive and negative error signals for the controller in the transient state by moving the actual potential value of the membrane above and below the resting potential value. The potential’s and the electric field’s magnitudes depend on the concentration and the geometry (the finite size of the membrane). The ions’ chemical nature comes into play only if the polarizability can differ for different molecules.
Figure 16: A block diagram of a PID controller in a feedback loop. r(t) is the desired process variable (PV) or setpoint (SP), and y(t) is the measured PV. (Wikipedia)
The overall control function of the general PID controller in Fig. 16 is described by
$u(t)=K_{p}\biggl(e(t)+\frac{1}{T_{i}}\int_{0}^{t}e(\tau)d\tau+T_{d}\frac{de(t)}{dt}\biggr)$ (46)
where $T_{i}=\frac{K_{p}}{K_{i}}$ is the time integration constant, $T_{d}=\frac{K_{d}}{K_{p}}$ is the derivative time constant. That is, in the world of a neuron, the "theory of everything" is
$V_{M}^{OUT}(t)=K_{p}\biggl(\overbrace{\underbrace{\overbrace{V_{M}^{IN}(t)}^{external}}_{clamping\ etc}+\overbrace{\underbrace{\frac{1}{T_{i}}\int_{0}^{t}V_{Wall}^{IN}(\tau)d\tau}_{membrane^{\prime}s\ wall}}^{integral,\ parallel\ RC}}^{Hodgkin-Huxley}+\overbrace{\underbrace{T_{d}\frac{dV_{M}^{IN}(t)}{dt}}}^{differential,\ serial\ RC}_{AIS}\,\biggr)$ (47)
Compare it to Eq. (40), which describes the serial $RC$ oscillator, and the usual equations describing HH’s hypothesis. The presence of some terms depends on the neuron’s actual stage; furthermore, the $R$ values in the parallel and serial $RC$ circuit differs. As discussed in section 4.1, subthreshold excitations provide direct experimental evidence that the description of the resting state includes both serial and parallel contributions.
The first term represents the constant effect of the external world (including synaptic inputs, clamping, and exciting brain tissue). The second term represents the resting state when no gradients are present. The voltage is mainly due to current through ion channels in the membrane’s wall; it represents a parallel $RC$ circuit. The third term represents the sum of the gradients due to all effects (the external world, the internal processes, and the outflow through the AIS); it represents a serial $RC$ circuit. HH considered only the first two terms plus the voltage gradient due to the condenser (but not the capacitive current). They missed that the input currents also produce a voltage gradient (clamping, per definitionem, supresses voltage gradients of the membrane) and that the output current through the AIS also produces a gradient. Furthermore, the equation explains that the external currents (say, step-functions’ rising and falling edges) also generate enormous changes in the output voltage. Similarly, an apparently "foreign" invasion (changing mechanically the location of charges by pressure, changing the concentration in one of the segments, ultrasound, magnetic pulse, and so on) generates a change in the concentration or position of charges on the neuronal membrane. This way, it generates a voltage gradient (as well as pressure gradient, furthermore, other changes), and so it may generate neuronal spikes.
From a control theoretical point of view, the condenser geometry and the ion concentrations set the ’always the same’ value (the set-point; see also Figure 1) and the currents serve to keep or restore that value. After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq. (19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, for discussing ’downhill’ channels, we must consider the charge-up of different components), the significant difference in current intensity needs a different discussion.
Given that the same number of charged ions must be present on the two sides of the membrane, their resultant surface density $\sigma_{A}$ must be the same. In the classic measurement by HH, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) is 380 and 390, respectively, while the concentration of the $Cl^{-}$ are 450 and 460; practically all equal, as our analysis has derived. The deviations are well within the uncertainty of the measured values. Notice that the effect is purely electrostatic, resulting in an asymmetric ion distribution; in that balanced state, no permeability is needed. If the membrane is (slightly) permeable, ions will move across the membrane until equilibrium is reached. The resulting potential difference is directly proportional to the concentration difference.
The classical condenser has an electric field, as the blue dashed line shows in Fig. 4. No field is outside the condenser or inside the condenser plates (see Fig. 3). There is a sudden jump in the field on the conductor’s surface (the armatures) and a constant field between the plates. (The charged layer can be ideally thin if electrons are sitting on the surface, but is of finite thickness on the opposite side where ions are on the surface. This latter effect is not discussed in classical electricity.)
The biological condenser behaves differently; see the red diagram line. An electrolyte (instead of an insulator) outside the condenser significantly alters the field’s structure. The electric field on the surface and between the plates changes by a factor of 3 to 4. Outside the plates, the polarization creates a field that changes logarithmically. An ion layer with finite thickness is built near the membrane’s surface. It is uniformly charged, so the field inside it is linear. At the internal surface of the condenser, it takes the value of the field calculated for the gap; on the other side of the layer, it takes the value of the logarithmic field at that position. The dashed line represents the gap field, the continuous line represents the charged layer’s field, and the dotted line represents the electrolyte’s field. We can observe the "correspondence principle": the discrete-continuous fields join seamlessly; furthermore, if the polarizability of the electrolyte decreases toward zero, the field value approaches the classical value.
The thickness of the charged layer has a significant impact on the resulting electric field (i.e., the neuron’s ’setpoint’). Suppose biological fragments are formed and settle on the membrane’s intracellular side due to the increased $\Delta z$. In that case, the resting potential (and the threshold potential) may change, potentially leading to neurological diseases. As [76] observed, the setpoint changes with temperature (the "thermodynamic electrical field" changes with the temperature, while the "electrical electrical field" does not, so their crossing point changes), reflected in the dynamics of the AP.
5.2 Robustness of the control circle
An exciting thought experiment is discussed in [5]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $Na^{+}]_{o}$, the outside or extracellular sodium concentration, is 145mM145\ mM, and $Na^{+}]_{i}$ is 12 mM. And suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential is set up to 66.6mV66.6\ mV. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In the absence of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law alone does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates its electric potential, which is a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electric potential. However, the thermal driving force exists, and the started ion transport quickly builds the electric layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage does not depend on its size, so it remains constant, and the voltage defined in this way adjusts the concentrations accordingly. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form an imperfect membrane with holes, establishing cellular electricity. As Fig. 9 depicts, the concentrations may adapt to the living conditions (whether living in seawater), but the general operating principles remain the same. The seawater environment enables the use of thinner membranes and lower membrane voltages; however, the same electric field is used in all cases.
Given that neurons share the external (global) concentrations (defined by the vast amount of ions in bulk), they provide a firmly fixed offset value to the always-the-same electric potential. This way, the operations of the individual neurons do not interfere. Furthermore, the internal (local) concentrations can adjust themselves even following a very rough charge perturbation while issuing an AP; moreover, when the living organism is growing.
5.3 Goldman-Hodgkin-Katz potential
It was already stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions adsorption on the membrane surface generates the membrane potential", for a review see [2]. The idea itself is nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability nor with ion absorption [20]. Furthermore, there is no idea in GHK about whether the ’setpoint’ (why that specific concentration or potential difference) is present and why the same potential is reset after rough perturbations such as issuing an AP or replicating a cell. The causality is reversed: the potential is a static concept and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of forming the gradients, mobility and permeability play a role, but not in determining the concentrations or the resting potential.
In our clear physical picture, the thermodynamic forces on the one side and on the other side of the membrane are summed. They counterbalance each other; furthermore, jointly the effect of the neuronal condenser, see Fig. 9. As we emphasized, the concentration gradients are ion-specific. Furthermore, as we discussed in connection with Eq.(2), the Nernst equation comprises a per-ion indefinite constant (a potential difference). To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is adding absolute concentrations on the different sides of the membrane, or changing the base of the logarithm used in a calculation to match the experimental value. It is not more than number magic. The potential is described by coupled equations as discussed in connection with Eq. (39).
The $Ca^{2+}$ ions do not fit into the GHK picture. One of the reasons why GHK cannot be good is that $Ca^{2+}]$ is omitted. Biologically, it is hard to believe that $Ca^{2+}$ does not participate in the game of life (especially since biology sees the need for $Ca^{2+}$ pumps). Physically, one concentration alone on one side of the membrane cannot keep balance; as the different ion exchange processes, Table 1 and Fig. 9 show. When adding a new ion to the solution, the sum concentration increases, and so the electrical force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$).
6 Summary
We provided a holistic, multi-level view of neuronal operation and described quantitatively the actual physical processes (instead of biophysics’ fairy tales about hypothesized lipid mechanisms); this way, we provided the promised ’new understanding’ (and the need for rewriting books in neuroscience). The classical neuron model, despite its ambition to be an electrical model, contradicts fundamental laws of electricity and conservation laws of science; furthermore, it cannot interpret other (thermodynamic, mechanical, optical) disciplinary phenomena. The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We proved that no compromise is needed: the pressure and the voltage observe the same physical process by different means. Neuroscience forgot that the neural current consists of ions, and they repel each other, in a confined volume.
In simple words, the neuron combines different disciplines; it works like a two-tact biological internal combustion engine, and can be described as a special Carnot-type thermodynamic engine. The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a vast pressure impulse, due to the electric repulsion of ions. The neuron’s volume remains practically unchanged, and the pressure wave moves the electrolyte (containing dissociated ions) out of the neuron through the only available output channel: the axon. In the first stroke, the pressure impulse invests energy into compressing the elastic membrane and starting a compression wave (soliton). In the second stroke, the elastic membrane pumps out electrolyte through the AIS, which "resists", and this way generates a potential wave, known as AP. The compression produces heat, observable as a temperature increase, and the expansion consumes heat, observable as a temperature decrease; these are well-understood processes in thermodynamics. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). In the discipline of electricity (when using the serial $RC$ model), it can be interpreted as a capacitive current, which in physiology is observed as hyperpolarization that inverts the polarity of the output voltage. The ions represent both mass and charge simultaneously, and correspondingly, in addition to the electric potential wave, mechanical, optical, and other feature changes can be observed. In different disciplines, pressure and voltage represent the same phenomenon, using concepts from various disciplines.
Research already provided the fundamental principles of a correct new ‘net electric’ neuron model, the laws of motion for describing biology, and provided initial steps toward a correct cross-disciplinary model (promising a unified and scientifically complete electric/thermodynamic model); all these open new perspectives for making exploratory experimental and theoretical research. The perspective includes an understanding of the actual thermodynamic and mechanical behavior of cells (and possibly, life) and the relationship between neural information and its carriers. The fact that the neuron receives through its synapses less than 3% of the energy it consumes during its AP cycle underscores that a neuron is a “summing amplifier” rather than a “repeater”. The time window is very narrow compared to the length of the operating cycle, so the neuron receives inputs only from a small subset of neurons connected to its synapses (the late input signals are gated, as Fig. 12 shows).
We have shown that the neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms and established that the neuron works as a simple control circuit where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentration of the cellular fluids define the nominal values of the electric and thermal gradients. Our results are in line with the conclusions "It is shown theoretically that the membrane potential enhancement becomes greater with decreasing the electrolyte concentration and pore radius" [19]. The ion currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By understanding that the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions) defines the value of the resting potential, we defied the decades-old fallacy that some mystic concerted operation of ion channels and some linear combination of purely-defined permeability values (the GHK equation) define the resting potential of neurons. By using the principles of classic electricity, we calculated, that neurons handle charge as electricity describes (electrolyte) condensers. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat absorption" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description.
We could theoretically underpin (see Fig. 9) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[1]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation on the membrane, as well as polarization in the electrolyte surrounding the membrane. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Considering the presence of electrolytes on the two sides of the membrane, a correct value for the resting potential can be derived numerically from the first principles. Those layers have a decisive impact on the neuron’s electrical operation. The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [8] of science must be considered.
By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [65] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Recognizing explicitly that biological systems use slow currents of charged ions instead of electrons solved some old mysteries and defied old fallacies.
Appendix A Speculations
By using the value of the force acting on a unit charge in the field across the membrane is
$F_{Na^{+}}=10^{7}*1.60217663*10^{-19}[V][C]=1.6*10^{-12}\ [N]$ (48)
we can estimate how the pressure of the neural cell increases due to the rush-in changes at the beginning of the AP. As evidence shows, the local potential at the internal surface of the membrane increases by $\Delta U=100\ [mV]$. This increase means a change in the force acting on an ion (see Eq. (48)) by $1.6*10^{-12}\ [N]$. When we assume $10^{7}$ rush-in ions and unchanged cell size, the total force acting on the membrane increases by $1.6*10^{-5}\ [N]$. This change in force means on the neuron’s $8*10^{-9}\ [m^{2}]$ surface (see [30], page 12) a pressure change
$\Delta P=\frac{1.6*10^{-5}\ [N]}{8*10^{-9}\ [m^{2}]}=2*10^{3}\quad\biggl[\frac{N}{m^{2}}\biggr]$ (49)
[…]
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed ([49], Fig. 3 quantitatively underpins the hypothesis) both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [49] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$ the speed at the exit of the ion channels can be estimated as $10^{4}~\frac{m}{s}$.
We can also estimate the work done when the rush-in charge extends the neuron, as the $\Delta P\times\Delta V$. We consider that $\Delta P$ remains constant, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $\Delta P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [70], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$. We can also estimate the energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [14]. To estimate the time of changing the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave) and it can be calculated as $F=m\frac{\Delta v}{\Delta t}$. Given that Eq. (48) provides the force and one must assume that the ion loses its $2*10^{5}~\frac{m}{s}$ speed, one arrives that it generates the shock wave in $\Delta t=5*10^{-8}~seconds$.
The consequences of the changes in the electric charge are large enough to explain why pressure wave and other mechanical changes [37] also start at the beginning of the AP, and other (such as optical, density) changes are accompanied by it. Nature invests energy also in the conventional way, a term $V\Delta P$, into the thermodynamics of neural operation; not only in the form of storing energy in the changed electric field. Thermodynamics and electricity, not to mention elasticity, must not be separated when discussing the neuronal energy business.
In line with the experimentally measured energy consumption and our calculation in section 4.5, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. That is, practically all the energy is invested as elastic energy, but the modulus is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in proximity to the membrane, but it creates a secondary, vast mechanical shock wave. By considering the electric repulsion force between the particles in the slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion consumes is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electric potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature (the temperature is more that two orders of magnitude higher, but the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect temperature change up to the range of fragments of millikelvins, depending on the measuring conditions). In the second phase of the AP, those more energetic ions that provide excess voltage (and so: excess local potential) above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value; another effect that can contribute to the general heat production and adsorption discussed above. In the original publication about solitons (essentially pressure waves) [78], a lot of measured and theoretical data are discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change; furthermore, that the electric energy is more than an order of magnitude smaller than the elastic energy (called soliton energy there).